### v1

# Data-Centric Foundation Models in Computational Healthcare: A Survey

Abstract.
The advent of foundation models (FMs) as an emerging suite of AI techniques has struck a wave of opportunities in computational healthcare. The interactive nature of these models, guided by pre-training data and human instructions, has ignited a data-centric AI paradigm that emphasizes better data characterization, quality, and scale. In healthcare AI, obtaining and processing high-quality clinical data records has been a longstanding challenge, ranging from data quantity, annotation, patient privacy, and ethics. In this survey, we investigate a wide range of data-centric approaches in the FM era (from model pre-training to inference) towards improving the healthcare workflow. We discuss key perspectives in AI security, assessment, and alignment with human values. Finally, we offer a promising outlook of FM-based analytics to enhance the performance of patient outcome and clinical workflow in the evolving landscape of healthcare and medicine. We provide an up-to-date list of healthcare-related foundation models and datasets at https://github.com/Yunkun-Zhang/Data-Centric-FM-Healthcare.
foundation models, large language models, data-centric AI, healthcare, AI alignment
copyright: noneccs: Applied computing Health informaticsccs: Computing methodologies Artificial intelligence
1. Introduction
The rise of foundation models (FMs) strikes a wave of breakthroughs for visual recognition (Radford et al., 2021; Kirillov et al., 2023; Ramesh et al., 2021), language understanding (Devlin et al., 2018; Brown et al., 2020; OpenAI, 2021, 2023), and knowledge discovery (Bommasani et al., 2021; Pan et al., 2023). In computational healthcare (Esteva et al., 2019; Aggarwal et al., 2021), FMs can handle a variety of clinical data with their appealing capabilities in logical reasoning and semantic understanding. Examples span fields in medical conversation (Shu et al., 2023; Yunxiang et al., 2023), patient health profiling (Christodoulou et al., 2019), and treatment planning (Nori et al., 2023). Moreover, given the strength in large-scale data processing, FMs offer a shifting paradigm to assess real-world clinical data in the healthcare workflow rapidly and effectively (Thirunavukarasu et al., 2023; Qiu et al., 2023).
FM research places a sharp focus on the data-centric perspective (Zha et al., 2023). First, FMs demonstrate the power of scale, where the enlarged model and data size permit FMs to capture vast amounts of information, thus increasing the pressing need of training data quantity (Vaswani et al., 2017). Second, FMs encourage homogenization (Bommasani et al., 2021) as evidenced by their extensive adaptability to downstream tasks. High-quality data for FM training thus becomes critical since it can impact the performance of both pre-trained FM and downstream models. Therefore, addressing key data challenges is progressively recognized as a research priority. In the healthcare system, collecting high-quality records could enable a comprehensive understanding of patient characteristics (imaging, genomics, and lab testing data) (Johnson and Khoshgoftaar, 2023; Singh, 2023; Alfasly et al., 2023). As illustrated, data-centric strategies promise to reshape clinical workflow (Rao et al., 2023; Juluru et al., 2021), enable precise diagnosis (Hunter et al., 2022), and uncover insights into treatment (Chen et al., 2023a).
Figure 1. Data-centric foundation models in computational healthcare.
Medical data challenges have posed persistent obstacles over decades, including multi-modality data fusion (Section 4), limited data volume (Section 5), annotation burden (Section 6), and the critical concern of patient privacy protection (Section 7) (Chen et al., 2021a; Hathaliya and Tanwar, 2020; Huang et al., 2020; Rajpurkar et al., 2022). To respond, the FM era opens up perspectives to advance data-focused AI analytics. Multi-modal FMs, as a concrete example, can offer scalable data fusion strategies for various data formats (Li et al., 2023a; Ding et al., 2023c). Meanwhile, the appealing trait of FM to generate high-quality data can greatly help address data quantity, scarcity, and privacy in the medicine and healthcare community (Tang et al., 2023a; Ding et al., 2023c; Chambon et al., 2022; Lu et al., 2022; Zhao and Bilen, 2023; Van Noorden and Perkel, 2023). To build responsible solutions for healthcare AI, the evolving perspective on AI-human alignment (Gabriel, 2020; Ngo et al., 2022) has become increasingly important. We discuss the necessity of the real-world applications of FMs aligned with human ethics, equity, and societal norms to reduce potential risks in performance assessment, ethical compliance, and patient safety (Liu et al., 2023d; Ozmen Garibay et al., 2023; Liang et al., 2022; Hathaliya and Tanwar, 2020). In the FM era, enabling AI-human alignment further underscores the significance of data focus, motivating us to prioritize the data-centric challenges in the landscape of computational healthcare.
[…]
Challenges [Modality Fusion
(Section 4) [Joint-Modal
Pre-Training [General FMs: CLIP (Radford et al., 2021); ALIGN (Jia et al., 2021); Florence (Yuan et al., 2021);
GLIP (Li et al., 2022); CoCa (Yu et al., 2022); FLAVA (Singh et al., 2022); CLAP (Elizalde et al., 2023)] [Medical FMs: ConVIRT (Zhang et al., 2022); BioViL (Boecking et al., 2022);
PLIP (Huang et al., 2023a); PubMedCLIP (Eslami et al., 2021); CheXzero (Tiu et al., 2022);
BiomedJourney (Gu et al., 2023)] ] [Leveraging LLMs [General FMs: Frozen (Tsimpoukelli et al., 2021); SimVLM (Wang et al., 2021);
Flamingo (Alayrac et al., 2022); BLIP-2 (Li et al., 2023c); LLaVA (Liu et al., 2023b);
NExT-GPT (Wu et al., 2023a)] [Medical FMs: LLaVA-Med (Li et al., 2023e); MedVInT (Zhang et al., 2023d);
Visual Med-Alpaca (Shu et al., 2023); Med-Flamingo (Moor et al., 2023b)] ] ] [Quantity
(Section 5) [Data Augmentation [Text: ChatGPT (Tang et al., 2023a)] [Imaging: Latent Diffusion (Packhäuser et al., 2022; Pinaya et al., 2022; Chambon et al., 2022);
Diffusion probabilistic model (Moghadam et al., 2023; Liu et al., 2022; Akrout et al., 2023);
[…]
CLIP (Zhang et al., 2023a); GLIP (Yi et al., 2023)] [Medical FM-based: REMEDIS (Azizi et al., 2023);
Biomedical LMs (Mishra et al., 2023; Zhang et al., 2023a)] ] [Internet-Scale
Data Curation [Datasets for healthcare FMs: PMC-15M (Zhang et al., 2023e);
Virchow (Vorontsov et al., 2023); OpenPath (Huang et al., 2023a); Quilt-1M (Ikezogwo et al., 2023);
PMC-OA (Lin et al., 2023b); HealthCareMagic-100k (Yunxiang et al., 2023)] ] ] [Annotation
(Section 6) [Text Annotation [Gilson et al. (Gilson et al., 2022); PathAsst (Sun et al., 2023); DeID-GPT (Liu et al., 2023e);
Med-PaLM 2 (Singhal et al., 2023); Tang et al. (Tang et al., 2023b)] ] [Image Annotation [Zero-shot SAM: SAM (Kirillov et al., 2023); Hu et al. (Hu and Li, 2023);
He et al. (He et al., 2023a); Mazurowski et al. (Mazurowski et al., 2023);
$\mathrm{SAM^{Med}}$ (Wang et al., 2023e)] [Adapting SAM: SAM features (Hu et al., 2023b); Poly-SAM (Li et al., 2023b);
COSMOS 553K (Huang et al., 2023b); Med SAM Adapter (Wu et al., 2023b);
SAMAug (Zhang et al., 2023f); MedSAM (Ma and Wang, 2023)] ] ] [Privacy
(Section 7) [Data Synthesizing [SoK (Hu et al., 2023a); Shibata et al. (Shibata et al., 2023); Tang et al. (Tang et al., 2023a);
Naresh et al. (Naresh et al., 2023); Dong et al. (Dong et al., 2022b); Sei et al. (Sei et al., 2024)] ] ] ]
Figure 2. An overview of healthcare data challenges and foundation model-based approaches mentioned in this survey paper.
In this survey, we offer a scoping perspective on developing, analyzing, and evaluating FM-focused approaches for healthcare. From a data-centric viewpoint as seen in Fig. 1, we emphasize the interplay between patients, healthcare data, and foundation models. We collect and discuss essential concepts, models, datasets, and tools for analyzing FMs (Fig. 2). Finally, we highlight emerging risks of applying FMs in healthcare and medicine regarding to privacy protection and ethical use. We offer promising directions for FM-based analytics to enhance the predictive performance of patient outcomes and streamline the clinical data workflows, ultimately leading to building better AI-human-aligned, data-focused tools, approaches, and systems in healthcare and medicine.
2. Foundation Models
Foundation models (FMs) are trained on the excessive-scale, wide-ranging data records towards high-level performance on downstream tasks (Bommasani et al., 2021). The key differentiation of general FM to classic deep learning models is on the scale of model size and training data. First, the success of FMs is built upon the Transformer-style model architecture (Vaswani et al., 2017), which can integrate large amounts of information through parallel computing and self-attention mechanisms (Zhou et al., 2023b). Second, FM training data contents normally encompass Internet-scale, multi-modal information with labeled and unlabeled annotations. With the information-rich pre-training, FMs exhibit a comprehensive understanding of concepts and their interrelationships, facilitating the application of knowledge to downstream tasks through adaptation.
As a concrete example of FM, large language models (LLMs) are pre-trained on Internet-scale corpora showing impressive semantic understanding and high-quality text generation (Brown et al., 2020; Touvron et al., 2023a; Chowdhery et al., 2022). In particular, LLMs emphasize the engineering design of input context, also known as the prompt. This design opens the door for human-and-machine interaction, allowing us to feed various prompt queries (e.g., text sequences) into a well-trained FM towards generating desired content or outcomes.
In this section, we discuss the essential concepts and useful capabilities in the cycle of general FM development, including large-scale pre-training (Section 2.1), fine-tuning (Section 2.2), and in-context learning (Section 2.3). These fundamental techniques are valuable to building a healthcare-focused FM.
2.1. Large-Scale Model Pre-Training
Large-scale model pre-training is an essential approach to building an FM from scratch. We discuss several core components including model size, data scale, and self-supervised learning methods, which play vital roles in developing pre-training techniques for building FMs.
2.1.1. Power of scale
The Transformer architecture serves as the backbone of FM, enabling efficient and scalable model training given its self-attention mechanism which allows parallel processing of input sequences (Vaswani et al., 2017). In the context of natural language processing (NLP), Transformer-based FMs have been trained on ever-larger corpora since 2018, demonstrating the power of scale (Zhou et al., 2023b). The document-level corpus enables the models to extract long-range information from long contiguous text sequences. For instance, Generative Pre-trained Transformer (GPT) (Radford et al., 2018) is a remarkable language model with 117M parameters pre-trained on more than 7,000 unique books (800M words) with various genres from the BooksCorpus dataset. Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) has 110M parameters for the base model and 340M for the large model. They include BooksCorpus and English Wikipedia (2,500M words) as the pre-training corpus.
With the increasingly massive amount of available training data, language model size has also increased (from millions to billions of parameters (Chowdhery et al., 2022)) to capture vast information embedded in the data. The example of Text-to-Text Transfer Transformer (T5) (Raffel et al., 2020) model has 11B parameters, pre-trained on the Colossal Clean Crawled Corpus (C4) (Raffel et al., 2020) dataset of approximately 750GB in size. GPT-3 (Brown et al., 2020) contains 175B parameters and is pre-trained on around 500B word tokens (i.e., fundamental units of text) from multiple sources. Large Language Model Meta AI (LLaMA) (Touvron et al., 2023a) with 7B to 65B parameters is pre-trained on public corpora containing trillions of tokens. Pathways Language Model (PaLM) (Chowdhery et al., 2022) pre-trained on 780B tokens has an increased size of 540B parameters.
[…]
2.1.2. Self-supervised learning
As the data scales up, the burden of human annotation for supervised model training becomes a practical challenge. Self-supervised learning (SSL) is widely used for FM pre-training without the need for labeled data (Devlin et al., 2018; Chen et al., 2021b; Radford et al., 2021). We categorize SSL research into two major approaches: input reconstruction and contrastive learning.
Input reconstruction is to minimize the distance between model-reconstructed data and the input data. For instance, BERT (Devlin et al., 2018) is pre-trained to perform masked language modeling, where a proportion of tokens are masked from the input sequence, and the model is purposely trained to predict the masked tokens. Similarly, masked image modeling is a common task for learning visual representations (He et al., 2022; Xie et al., 2022), where image patches are masked during model training and then the model is required to reconstruct the masked patches. Another example is GPT (Radford et al., 2018) which is pre-trained to predict the next token of a corpus given the previous tokens in the text sequence.
Contrastive learning is to maximize the similarity of similar data samples (positive samples) in the embedding space while minimizing that of dissimilar data samples (negative samples) (Chen et al., 2020). SimCLR (Chen et al., 2020) and MoCo-v3 (Chen et al., 2021b) are examples of applying two random data augmentations to each training image to generate positive samples and recognize different images as negative samples to train a vision transformer. Besides, Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) views image-text pairs as positive samples and irrelative image and text as negative samples to learn aligned visual and text representations.
Growing approaches combine input reconstruction and contrastive learning for more robust representations. For instance, DINOv2 (Oquab et al., 2023) trains a student network to match image representations with a slowly updating teacher network in a contrastive manner, where the two networks take different crops of the same image as input. The student network is required to reconstruct masked image patches. DINOv2 representations achieve outstanding performance on most visual benchmarks at both image and pixel levels. In healthcare, studies have shown that pre-training large models using self-supervised learning on large-scale medical unlabeled data can also achieve effective and efficient adaptation on downstream medical tasks, reducing annotation burden (Krishnan et al., 2022). For example, PubMedBERT (Gu et al., 2021) is a self-supervised BERT model pre-trained on 14M PubMed abstracts from scratch without manual annotation, achieving state-of-the-art results on various biomedical NLP tasks. MoCo-CXR (Sowrirajan et al., 2021) pre-trains a visual model on abundant unlabeled chest X-ray images, resulting in robust image representations that exceed supervised learning performance.
Figure 3. Foundation model (FM) in healthcare.
2.2. Fine-Tuning
The paradigm of pre-training and fine-tuning has dominated deep learning since self-supervised learning enabled large-scale model pre-training. In detail, fine-tuning refers to updating pre-trained model parameters through forward and backward passes and gradient descent on task-specific supervised data. Fine-tuning is a critical technique in constructing medical domain-specific FMs (Lee et al., 2020; Gema et al., 2023; Li et al., 2023e). However, conventional fine-tuning poses significant challenges including high cost, extensive data requirements, and limited generalization to unseen tasks. By contrast, we introduce two essential techniques beyond basic fine-tuning to address those issues: parameter-efficient fine-tuning and instruction tuning.
2.2.1. Parameter-efficient fine-tuning
Parameter-efficient fine-tuning (PEFT) is a family of fine-tuning approaches that only updates a small set of model parameters while keeping most of the pre-trained weights fixed. These methods alleviate critical challenges caused by updating all parameters of a large FM during full fine-tuning, such as the excessively high cost, overfitting on small downstream datasets, and the risk of catastrophic forgetting (Ding et al., 2023a). PEFT approaches excel in scenarios with limited data availability and effectively retain the valuable knowledge acquired through FM pre-training. We discuss essential PEFT methods. BitFit (Zaken et al., 2021) fine-tunes only the bias parameters of the linear layers in the model. Adapter (Houlsby et al., 2019) inserts lightweight modules into the transformer blocks. Prompt tuning (Lester et al., 2021) prepends trainable tokens to the tokenized input sequence of the first transformer layer. Prefix tuning (Li and Liang, 2021) is like prompt tuning, except that trainable tokens are prepended to the input of all transformer layers. Low-Rank Adaptation (LoRA) (Hu et al., 2021) reparameterizes the weight matrices with multiplications of low-rank matrices. PEFT methods demonstrate potential effectiveness in healthcare (Gema et al., 2023).
2.2.2. Instruction tuning
Instruction is defined as the linguistic description of a task along with its corresponding task-specific data sample. Instruction tuning refers to fine-tuning FMs on supervised instruction datasets with LLMs helping to understand the instruction (Ouyang et al., 2022). This method enhances zero-shot performance on new tasks and improves the generalization capability of the fine-tuned FM. The outcome of instruction tuning is influenced by both the proficiency of the pre-trained FM and the quality of the instruction-following data (Taori et al., 2023). For instance, Fine-tuned LAnguage Net (FLAN) (Wei et al., 2021) is an LLM fine-tuned on tens of NLP datasets via natural language instructions, outperforming GPT-3 on most of the zero-shot evaluation datasets. In addition, InstructGPT (Ouyang et al., 2022) and ChatGPT (OpenAI, 2021) are LLMs fine-tuned on instructions with human feedback to promote user alignment, resulting in chatbots with more truthful outputs. Alpaca (Taori et al., 2023) is a LLaMA model fine-tuned on 52K InstructGPT-generated instructions, achieving performance similar to InstructGPT but with a significantly smaller model size. Instruction tuning has also been applied to build healthcare FMs (Li et al., 2023e; Yunxiang et al., 2023).
2.3. In-Context Learning
Large language models (LLMs), as a concrete example of FM, bring a promising learning paradigm termed in-context learning (ICL), also known as prompt engineering (Brown et al., 2020). Prompt essentially represents an input query that guides the model’s output that has proven to greatly impact the model’s performance even without the need for fine-tuning. By merely adjusting the input query of LLMs during inference instead of updating any parameters of LLMs, we can flexibly prompt LLMs to generate desired outputs without fine-tuning them on downstream tasks (Brown et al., 2020; Dong et al., 2022a).
ICL promises to extend powerful LLMs into downstream tasks by injecting contextual information. We introduce essential ICL strategies related to FM applications. Zero-shot is the simplest type of prompt, where we directly assign a task to the model, e.g., “Tell me a joke”. Few-shot gives the model a few demonstrations of the task. For example, “bird = oiseau. cat = chat. dog =”. In this case, we prompt the model to translate the English word “dog” to French by giving two examples. Chain-of-thought (CoT) (Kojima et al., 2022) prompts LLMs to generate the reasoning path (i.e., step-by-step thoughts) for addressing complicated problems. For example, “Q: I had 10 apples. I gave my neighbor 3 apples. Then I bought 4 apples and ate 2. How many apples do I have now? A: Let’s think step by step:”. Based on CoT, self-consistency (Wang et al., 2022b) aggregates a diverse set of reasoning paths to get the most consistent answer. Recently, tree-of-thought (ToT) (Yao et al., 2023) further builds a tree of reasoning paths and applies tree search methods to obtain the best path. Self-refine (Madaan et al., 2023) prompts an LLM to provide feedback for its own outputs and refine itself. We can even prompt an LLM to write prompts.
FMs start to demonstrate potential generalization power in the medical field using ICL techniques (Roy et al., 2023; Lyu et al., 2023). For instance, ChatGPT can effectively translate radiology reports into plain language by using well-designed prompts that inform the model about the structure of the report and contents of each paragraph (Lyu et al., 2023). Also, the general-purpose SAM performs well on abdominal CT organ segmentation when provided point and bounding box prompts in an oracle manner (Roy et al., 2023).
3. Foundation Models in Healthcare
The growth of FM analytics offers insights into healthcare applications (Qiu et al., 2023; Wornow et al., 2023; Zhang and Metaxas, 2023). We review key techniques, tools, and applications addressing multiple aspects of FM in healthcare. We exhibit how general-purpose FMs can be applied in the healthcare field (Section 3.1). We present medical-focused FMs and demonstrate pre-training benefits gained from general FMs (Section 3.2).
3.1. Adapting General Foundation Models in Medicine and Healthcare
Research efforts have started to assess FM’s superior capability in the medical domain (Nori et al., 2023; Gema et al., 2023; Roy et al., 2023). In these studies, we identify two core techniques including parameter-efficient fine-tuning (PEFT) and in-context learning (ICL).
3.1.1. Adapting via parameter-efficient fine-tuning (PEFT)
PEFT methods have been applied to adapt FMs to medical tasks. For instance, Dutt et al. (Dutt et al., 2023) demonstrate that PEFT methods significantly outperform full fine-tuning of FMs in data-limited scenarios for medical image classification and text-to-image generation tasks. Gema et al. (Gema et al., 2023) propose a two-stage PEFT framework to adapt LLaMA (Touvron et al., 2023a) to a broad range of clinical tasks. In this work, the first stage applies LoRA (Hu et al., 2021) to fine-tune LLaMA on clinical notes, building Clinical LLaMA-LoRA, a clinical FM; the second stage again applies LoRA to adapt the clinical FM to downstream tasks. They also demonstrate that LoRA, among major choices of PEFT methods, works ideally for the clinical domain adaptation. Similarly, Van Veen et al. (Van Veen et al., 2023) apply LoRA to fine-tune T5 models (Raffel et al., 2020; Lehman and Johnson, 2023) for radiology report summarization. They also apply LoRA together with in-context learning for clinical text summarization tasks, showing improved performance over human experts (Veen et al., 2023).
3.1.2. Adapting via in-context learning (ICL)
ICL has proven to be effective in adapting FMs, especially large language models (LLMs), to a variety of healthcare tasks. With carefully designed task-specific input context (i.e., prompts), the FM can perform well on healthcare tasks without modifying any model parameters. For instance, Nori et al. (Nori et al., 2023) evaluate GPT-4 (OpenAI, 2023) on the United States Medical Licensing Examination (USMLE) without specially crafted prompts. GPT-4 shows its promising zero-shot performance without adding relevant medical context data. Lyu et al. (Lyu et al., 2023) leverage ChatGPT (OpenAI, 2021) to translate radiology reports into plain language for report understanding and translation. The experiments show that by using a clearer and more structured prompt, the overall translation quality can be increased. Roy et al. (Roy et al., 2023) show that SAM (Kirillov et al., 2023) has a great generalization ability in abdominal CT organ segmentation, through its point/bounding box prompting. Deng et al. (Deng et al., 2023) evaluate the zero-shot performance of SAM on tumor segmentation, non-tumor tissue segmentation, and cell nuclei segmentation on whole slide images (WSI), demonstrating that SAM performs well on large connected objects on pathological scans. Chen et al. propose Diagnosis of Thought (DoT) prompting (Chen et al., 2023b) to assist professionals with cognitive distortion detection. DoT diagnoses mental illness by prompting LLMs to sequentially perform subjectivity assessment, contrastive reasoning, and schema analysis.
3.2. Pre-Training Healthcare Foundation Models
Researchers make efforts to pre-train FMs based on large-scale unlabeled healthcare data for health record examination (Gu et al., 2021; Alsentzer et al., 2019; Singhal et al., 2022), medical imaging diagnosis (Azizi et al., 2023; Wang et al., 2022c), and protein sequence analysis (Lin et al., 2023a; Cheng et al., 2023a). In principle, the pre-training process can be summarized into two major aspects: pre-training strategy and model initialization.
3.2.1. Pre-training strategy
Healthcare FM pre-training typically utilizes a range of pre-training strategies derived from general-domain FMs due to their potential generalization power.
The first pre-training strategy is masked language/image modeling, following BERT (Devlin et al., 2018) and masked autoencoder (MAE) (He et al., 2022). For instance, SciBERT (Beltagy et al., 2019) and PubMedBERT (Gu et al., 2021) are pre-trained on multi-domain scientific publications and biomedical domain-specific corpora respectively, based on BERT strategy. BioLinkBERT (Yasunaga et al., 2022) utilizes links between biomedical documents and is pre-trained with masked language modeling and document relation prediction tasks. BioGPT (Luo et al., 2022) is pre-trained on PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/ abstracts following GPT-2 (Radford et al., 2019) for generative language tasks. RETFound (Zhou et al., 2023a) is an FM for retinal image disease detection, pre-trained on a large collection of unannotated retinal images to reconstruct input images with 75% masked patches, following MAE. Similarly, General Expression Transformer (GET) (Fu et al., 2023) is an FM for modeling transcriptional regulation across 213 human cell types. GET is pre-trained to predict the motif binding scores of masked regulatory elements in the input to learn regulatory patterns.
Contrastive learning is another important pre-training strategy for medical FMs. For example, REMEDIS (Azizi et al., 2023) is a medical vision model pre-trained via contrastive learning to extract representative visual features for medical images. Vision-language models such as MedCLIP (Wang et al., 2022c), MI-Zero (Lu et al., 2023b), and PLIP (Huang et al., 2023a), are contrastively pre-trained on domain-specific image-text pairs. They achieve positive performance on zero-shot image classification tasks in radiology and pathology.
Beyond these pre-training strategies, inspiration has been drawn from language models’ ability to process sequential inputs, leading to large pre-trained models for protein sequence tasks. ESM-2 (Lin et al., 2023a) is a Transformer model with 15B parameters pre-trained on millions of protein sequences to predict protein structure directly from amino acid sequences, fast and accurately. ESM-2 illustrates the immense potential of LLMs to learn patterns in protein sequences across evolution. AlphaMissense (Cheng et al., 2023a) pre-trains an AlphaFold-like model (Jumper et al., 2021) to predict protein structure via protein language modeling. It then fine-tunes the model with an additional variant pathogenicity classification objective on human and primate variant population frequency databases. AlphaMissense achieves state-of-the-art performance on missense variant pathogenicity prediction.
3.2.2. Model initialization
Healthcare FM pre-training benefits from utilizing general FMs as the initial model to leverage their massive information. With proper model initialization, we recognize that much less data and fewer training epochs are needed for domain-specific pre-training of medical FMs. For example, BioBERT (Lee et al., 2020) is pre-trained on PubMed abstracts and PubMed Central (PMC [^2] https://www.ncbi.nlm.nih.gov/pmc/) full-text articles with BERT initialization and demonstrates better performance compared with BERT and previous state-of-the-art models when fine-tuned on three biomedical language tasks. Similarly, BlueBERT (Peng et al., 2019) continues BERT pre-training on PubMed abstracts and MIMIC-III clinical notes (Johnson et al., 2016). Alsentzer et al. (Alsentzer et al., 2019) build clinical text models initialized from BERT and BioBERT, and show a significant improvement on five clinical language tasks. PMC-LLaMA (Wu et al., 2023d) is an open-source language model fine-tuned from LLaMA (Touvron et al., 2023a) on biomedical academic papers. Singhal et al.  (Singhal et al., 2022) apply instruction tuning on Flan-PaLM (Chung et al., 2022) to obtain Med-PaLM model, achieving state-of-the-art performance on a broad range of medical question answering (MedQA) tasks. PubMedCLIP (Eslami et al., 2021), a CLIP model fine-tuned on medical image-text pairs from PubMed articles, shows outstanding performance on medical visual question answering (MedVQA) tasks. However, medical data heterogeneity has long been an obstacle to FM pre-training, generalization, and evaluation. As we see that medical data records reflect the complex nature of human disease complexity in clinics, overcoming this data heterogeneous challenge is crucial to enable FM robustness in real-world healthcare applications (Hathaliya and Tanwar, 2020; Chen et al., 2021a; Rajpurkar et al., 2022; Liang et al., 2022; Huang et al., 2020).
4. Multi-Modal Data Fusion
Data fusion is a useful strategy to aggregating the information of various medical data modalities towards improved decision-making in healthcare. Standard fusion methods can be conceptually categorized into three types (Stahlschmidt et al., 2022; Huang et al., 2020): early fusion, joint fusion, and late fusion. In early fusion, data from different modalities are combined at the input and passed to the subsequent network (Park et al., 2020; Xie et al., 2019; Zhao et al., 2021; Suresh et al., 2017; Bichindaritz et al., 2021; Fu et al., 2020). Joint fusion processes data from each modality through independent networks before fusing their representative feature maps as the input of the subsequent network (Braman et al., 2021; Spasov et al., 2018; Kawahara et al., 2018). Late fusion processes data from each modality through individual networks and fuses their output to make the final outcomes (Reda et al., 2018; Qiu et al., 2018). However, conventional methods suffer from a lack of scalability, generalizability, and cross-modal understanding (Huang et al., 2020). For example, early and joint fusion methods often require sufficient task-specific training data and computational resources, while late fusion methods usually show weaknesses in multi-modal information integration (Budikova et al., 2017). Figure 4 illustrates the healthcare data modalities, data fusion strategies, and associated healthcare tasks.
We recognize that multi-modal FMs can enable a more scalable, generalizable, and comprehensive data fusion (Ramesh et al., 2021; Radford et al., 2021; Bao et al., 2022; Elizalde et al., 2023; Li et al., 2023a). We thus discuss the benefits of multi-modal FMs from two primary aspects. First, joint-modal pre-training of multi-modal FMs enhances data fusion to a massive scale with transferability to downstream healthcare tasks. Second, large language models (LLMs) possess strong comprehension and reasoning abilities, which can be leveraged to understand cross-modality interaction given the aggregated multi-modal inputs.
Figure 4. Multi-modal fusion of healthcare data in the FM era. Conventional fusion approaches are enhanced by joint-modal pre-training and comprehensive FMs such as LLMs, enabling downstream applications such as medical QA, drug discovery, and diagnosis.
4.1. Data Fusing via Joint-Modal Pre-Training
FMs can handle multiple modalities via pre-training on massive-scale paired multi-modal data in a joint-modal mode to obtain a high-level understanding of inter-modality relationships. For instance, CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), and Florence (Yuan et al., 2021) adopt image-caption pairs collected from the Internet. These models extract image and text features with separated networks and project both features into a shared latent space, enabling a joint training network. These models have shown great zero-shot transferability when testing on unseen image-text combinations, including medical images and healthcare reports (Liu et al., 2023f). In addition, GLIP (Li et al., 2022), CoCa (Yu et al., 2022), and FLAVA (Singh et al., 2022) introduce more advanced fusion modules, such as cross-attention and generative models, to better explore cross-modality interaction during pre-training. Besides vision-language models, there exist models for other modalities with joint-modal pre-training. For example, CLAP (Elizalde et al., 2023) is an audio-language model pre-trained on audio-caption pairs.
Joint-modal pre-training has been applied for healthcare FMs, especially in radiology, where chest X-ray images and reports are frequently paired. ConVIRT (Zhang et al., 2022) is pre-trained on numerous pairs of chest X-rays and radiology reports, fusing vision and language modalities via contrastive learning. BioViL (Boecking et al., 2022) trains a visual model together with an in-domain (CXR) language model by optimizing contrastive loss and masked language modeling. In addition, PubMedCLIP (Eslami et al., 2021), CheXzero (Tiu et al., 2022), and PLIP (Huang et al., 2023a) extend the CLIP model to image-text data in biomedicine, radiology, and pathology respectively. BiomedJourney (Gu et al., 2023) is a latent diffusion model (Rombach et al., 2022) trained on tuples of pre- and post-disease progression images from patients coupled with a GPT-generated natural language description of the progression, enabling the generation of counterfactual medical images.
4.2. Data Fusing via LLMs
Transformer-style LLMs possess powerful semantic understanding capability via the attention mechanism (Brown et al., 2020), which can be transferred to multi-modal settings. To be specific, data from different modalities can be aggregated as the prompt input of an LLM (i.e., a sequence of tokens). These combined multi-modal inputs are then fused through the Transformer blocks in the LLM, exchanging information via attention layers (Wang et al., 2023a; Liu et al., 2023a, b). For instance, Frozen (Tsimpoukelli et al., 2021) concatenates image features and text tokens as the input of a pre-trained generative language model. This design enables the vision encoder to update based on information from both modalities. Similarly, SimVLM (Wang et al., 2021) fuses image features and text tokens with BERT-like bi-directional attention layers and trains the whole model together. Flamingo (Alayrac et al., 2022) takes as input visual data interleaved with text and inserts gated xattn-dense layers into the language model to condition it on visual inputs. BLIP-2 (Li et al., 2023c) adopts a well-designed Q-Former to bridge the gap between pre-trained vision and language models before passing into the LLM. In addition, LLaVA (Liu et al., 2023b) feeds image features and language instructions into LLaMA to train a multi-modal chatbot. NExT-GPT (Wu et al., 2023a) is a unified agent understanding multiple modalities, including text, images, videos, and audio.
Since LLMs show great potential in the medical field (Nori et al., 2023), these methods are applied to medical data fusion as well. Med-Flamingo (Moor et al., 2023b) is an open-ended MedVQA model based on Flamingo, pre-trained on paired and interleaved medical image-text data. MedVInT (Zhang et al., 2023d) applies instruction tuning to a LLaVA-like multi-modal model with MedVQA data to pre-train a generalizable MedVQA foundation model. Visual Med-Alpaca (Shu et al., 2023) is designed to handle multi-modal biomedical tasks, based on the LLaMa-7B architecture and trained with an instruction set designed cooperatively by GPT-3.5-Turbo and human experts. Moreover, LLaVA-Med (Li et al., 2023e) curates biomedical instruction-tuning data by prompting language-only GPT-4 to generate multi-round questions and answers about biomedical images. Surprisingly, GPT-4 generates high-quality visual question-answering conversations even if it has access only to the text. These models show outstanding zero-shot/few-shot performance on downstream medical tasks that require multi-modality understanding.
Current multi-modal FMs predominantly focus on language and vision, given their ubiquitous use as the most common modalities. For FM applications in the healthcare field, there are only handful examples including genomics (Ding et al., 2023b), electrocardiogram (Chang et al., 2023a), electroencephalogram (Cui et al., 2023a), and tabular data (Wang et al., 2023c). We believe that joint-modal pre-training techniques, combined with LLMs’ logical reasoning capabilities, could benefit healthcare data fusion of modalities beyond the scope of language and vision.
5. Data Quantity
Towards applying FMs in the healthcare system safely and responsibly, we must address a variety of data quantity hurdles. The trade-off of data quantity is between the limited information provided by a constrained dataset and the essential information required for training a robust model on the specific healthcare task. Such a disparity can result in inadequate downstream model training, yielding non-robust, inaccurate, and even unreliable model outcomes. However, public patient data records are often scarce due to the rigorous patient privacy protection regulations. In addition, we must deal with the costly real-world dataset curation processes including data collection, cleaning, and annotation. While deep learning has shown its promise, the barrier of limited data access in healthcare continues to be a significant roadblock for scalable AI-based research.
The emergence of FMs holds the potential to alleviate the data quantity challenges in healthcare applications. The key design of FMs is to pre-train models on large-scale datasets that consist of billions of samples, obtaining a vast volume of information to compensate for the limited data scenarios in downstream healthcare tasks. In this section, we discuss representative approaches to handling data quantity challenges using FMs from the perspectives of data augmentation (Section 5.1) and data efficiency (Section 5.2). We then introduce the curation of large-scale healthcare datasets from the Internet to support healthcare FM pre-training (Section 5.3).
5.1. Data Augmentation
Data augmentation is a common strategy in machine learning that has proven to be effective in addressing the issue of data sample limitation (Shorten and Khoshgoftaar, 2019; Shorten et al., 2021). Conventional augmentation techniques include resizing, clipping, and flipping for images (Krizhevsky et al., 2012; Szegedy et al., 2016; DeVries and Taylor, 2017; Cubuk et al., 2018) and synonym replacement, random insertion, and back translation for text (Sennrich et al., 2015; Wei and Zou, 2019; Xie et al., 2020; Edunov et al., 2018). Data augmentation has also been applied within the healthcare domain (Kang et al., 2023; Tellez et al., 2019; Shen et al., 2022). Yet, these techniques only manipulate the existing data samples and maintain limited information entropy since no external information is introduced beyond the existing data distribution itself.
Information-rich generative FMs have brought a remarkable shift to healthcare data augmentation for both imaging and text (Kazerouni et al., 2023; Nori et al., 2023). Pre-trained with a vast set of knowledge, FMs enable the transfer of general insights to the healthcare domain beyond the scope of limited datasets.
[…]
Meanwhile, FMs demonstrate positive performance in clinical text mining through textual data augmentation. Employing ChatGPT (OpenAI, 2021) to generate high-quality synthetic data with labels has proved beneficial for fine-tuning a local model for a downstream task (Tang et al., 2023a). This approach effectively addresses the challenge of data quantity in clinical text mining. By utilizing ChatGPT to generate synthetic data, the burden of extensive data collection is significantly reduced.
5.2. Data Efficiency
Data-efficient approaches aim to improve downstream data efficiency with the support of FMs to reduce the data volume required by downstream tasks. Serving as a bridge to connect massive upstream data and limited downstream data, these methods help improve data efficiency and alleviate the data quantity challenges in healthcare. Research efforts have shown that by incorporating knowledge from the pre-trained general-domain FMs, healthcare datasets with limited sizes can yield satisfactory results. For example, CITE (Zhang et al., 2023a) explores the adaptation of general vision FMs, such as CLIP (Radford et al., 2021) and INTERN (Shao et al., 2021), to comprehend pathological images, shedding light on the utilization of medical domain-specific text knowledge to enhance data-efficient pathological image classification. Wu et al. (Wu et al., 2023c) demonstrate that general large language models (LLMs) such as ChatGPT (OpenAI, 2021) and GPT-4 (OpenAI, 2023) exhibit strong capabilities in handling radiology natural language inference (NLI) tasks even with limited data.
In addition to general FMs, leveraging healthcare-focused FMs represents a reasonable direction for pursuing data efficiency, significantly reducing the data requirement in downstream healthcare applications. To illustrate, REMEDIS (Azizi et al., 2023), a data-efficient training strategy combining large-scale supervised transfer learning with self-supervised learning, demands minimal task-specific customization. REMEDIS requires only 1–33% of downstream data for fine-tuning to match the performance of supervised models retrained using all available data in out-of-distribution settings. As for medical text data, Mishra et al. (Mishra et al., 2023) demonstrate that pre-trained medical text encoders exhibit promising performance, particularly in handling low-prevalence diseases. This finding holds promise for addressing data quantity challenges in other healthcare domains. Moreover, Yi et al. (Yi et al., 2023) introduce continual learning, including sequential learning and rehearsal learning, based on medical FMs as a practical and data-efficient learning paradigm, offering a promising avenue to tackle data quantity challenges effectively.
5.3. Internet-scale Data Curation
The data quantity challenge in the healthcare system strongly limits the performance of downstream applications. FM provides a novel means to utilize large-scale public data collected from the Internet. To support healthcare FM pre-training, the curation and process of Internet-scale healthcare datasets is critical. Ensuring representative collection of healthcare data requires high-quality data sources and effective data extraction strategies. One common approach is to extract text and images from PubMed Central (PMC), an Internet-scale archive of biomedical journal literature, or other well-established online medical databases. For instance, PMC-OA (Lin et al., 2023b) and PMC-15M (Zhang et al., 2023e) extract image-caption pairs from PMC articles. Virchow (Vorontsov et al., 2023) collects pathology scans from the Memorial Sloan Kettering Cancer Center (MSK), a cancer treatment and research institution. Other examples leverage large Internet platforms. OpenPath (Huang et al., 2023a) utilizes pathology hashtags to collect tweets as well as their replies containing certain keywords or with high numbers of likes. Quilt-1M (Ikezogwo et al., 2023) searches the YouTube platform for histopathology videos and obtains corresponding text through speech-to-text techniques and LLM postprocessing. HealthCareMagic-100k (Yunxiang et al., 2023) collects patient-physician dialogues from an online medical consultation platform.
Figure 5. Foundation models address data quantity and data annotation challenges. Left: Foundation models can mitigate data quantity limitation by data augmentation and improved data efficiency. Right: Foundation models can help both healthcare text and medical image annotation.
6. Data Annotation
Data annotation is a critical step in computational healthcare, involving the labeling of medical data, such as medical images, electronic health records, or genomic sequences (Belle et al., 2015). By assigning informative metadata or class labels, data annotation enriches the raw dataset with a nuanced layer of human expertise and contextual understanding, providing valuable insights for healthcare education, diagnostics, and artificial intelligence applications (Lee et al., 2017; Willemink et al., 2020). As such, data annotation is pivotal in the ongoing quest to harness the power of FM in diagnosing diseases, personalizing treatment plans, and ultimately paving the way for a more informed and efficient healthcare system (Razzak et al., 2018). However, numerous challenges in data annotation remain to be addressed, including the shortage of professional annotators and the complexity of the annotation process (Dgani et al., 2018).
The scalability of FMs allows us to alleviate the labor-intensive healthcare data annotation. Especially, FM-based scalable annotation paradigm is developing rapidly in the medical field for the following reasons. First, FM core techniques such as tokenization and modality fusion ability intend to erase the boundaries of different data modalities. Second, vast volumes of data are accumulating rapidly, paving the path for the development of a potent world model that comprehends more modalities as special languages (Wang et al., 2022a). In this section, we demonstrate that FM provides a promising avenue to simplify data annotation in healthcare. In the context of healthcare text annotation, large language models (LLMs) have shown potential capability on clinical language tasks including medical question answering (Section 6.1). As for medical imaging, using multi-modal FMs and image segmentation FMs promises to improve the efficiency of excessive image captioning and annotation (Section 6.2).
6.1. Healthcare Text Annotation
Healthcare text annotation is a process that extracts and categorizes critical information within a wide range of healthcare texts. This process plays a crucial role in the healthcare system to enhance data quality and knowledge discovery (Xia and Yetisgen-Yildiz, 2012). However, healthcare text annotation is labor-intensive and time-consuming, observed from studies in medical condition labeling (Briesacher et al., 2008), patient record annotation (Hogan and Wagner, 1997), and key medical entity identification (Abacha and Zweigenbaum, 2011). To streamline a text annotation process, an emerging direction is to harness general-purpose LLMs. For instance, Gilson et al. (Gilson et al., 2022) reveal that ChatGPT (OpenAI, 2021) achieves comparable human-level performance on medical question-answering tasks, showcasing the potential for LLMs to enhance the efficiency of text annotation in a more flexible manner. Med-PaLM 2 (Singhal et al., 2023), an LLM fine-tuned with healthcare data, has performed comparably to human clinical professionals on medical knowledge retrieval, reasoning, and question answering. This suggests that integrating Med-PaLM 2 into medical data annotation workflows could lead to more accurate and efficient annotation processes, harnessing the power of LLMs in the medical domain.
While LLMs have showcased the ability to produce grammatically accurate and human-like text (Thirunavukarasu et al., 2023; Gilson et al., 2022), they still encounter challenges in specialized healthcare contexts. A noticeable knowledge gap in healthcare text still remains between general-domain FMs and medicine professionals, resulting in inaccurate responses to realistic patient inquiries, such as in cardiovascular disease prevention (Sarraju et al., 2023). Besides, Liao et al. (Liao et al., 2023) also highlight that the linguistic attributes of medical text generated by ChatGPT considerably differ from those of human experts.
Advancements are required to improve the utilization of text-based FMs for improving the performance of medical text annotation. PathAsst (Sun et al., 2023) leverages FMs as a generative AI assistant to transform predictive analytics in pathology. Using ChatGPT and GPT-4, PathAsst generates over 180,000 instruction-following samples to invoke pathology-specific models and facilitate effective interactions based on input images and user inputs. Remarkably, generative FMs have shown the potential of text annotation to assist and enhance pathology diagnosis. Also, Liu et al. (Liu et al., 2023e) develop DeID-GPT as a de-identification framework powered by GPT-4 (OpenAI, 2023), designed to detect and eliminate identification information from clinical text records automatically. In contrast to conventional methods, DeID-GPT utilizes in-context learning, which requires no code or procedural changes when applying the solution to different hospitals, different languages, and different data formats. It improves reliability by concealing sensitive information while preserving the original structure of the clinical text. Moreover, Tang et al. (Tang et al., 2023b) utilize the language model BART (Lewis et al., 2019) to enhance medical dialogue generation by focusing on domain-specific terminology. Empirical results highlight the model’s efficiency in generating responses based on the historical context of dialogue exchanges between medical professionals and patients.
6.2. Medical Image Annotation
Medical image annotation includes outlining and labeling anatomical structures, such as organs, tumors, blood vessels, or bones in histopathology and radiological images (Gao et al., 2021, 2022). Such a task can also involve annotating regions of interest in histopathological slides to identify cancerous cells or specific tissue types of clinical significance. Medical image annotation plays a pivotal role in modern healthcare to help extract and interpret valuable information from complex medical image examinations. This process is essential for both clinical diagnosis and research purposes, as healthcare professionals require computer-aided algorithms to efficiently identify, search, manage, and interpret disease indicators (Yin and Bi, 2018).
Robust segmentation models can be transformed into powerful tools for healthcare image annotation, thus leading to a substantial reduction in annotation costs and the facilitation of large-scale dataset curation spanning a wider range of healthcare imaging applications (Gao et al., 2023). For instance, Qu et al. (Qu et al., 2023) introduce an AI-driven systematic methodology that accelerates organ segmentation annotation by 500 times. While there have been advances in medical image annotation using FMs, the absence of a consensus on state-of-the-art benchmarks remains a significant challenge.
The advent of segmentation FMs, such as the Segment Anything Model (SAM) (Kirillov et al., 2023), has made promising progress on automated image segmentation and annotation. By training with over 1 billion masks sourced from 11 million natural images, SAM can execute zero-shot image segmentation using diverse input prompts, including masks, boxes, and points information. The potential of harnessing SAM for medical annotation emerges as a promising direction for analyzing complex medical objects. Numerous studies have applied FM-based medical segmentation to gain insightful results. Wang et al. (Wang et al., 2023e) present $\mathrm{SAM^{Med}}$, an enriched framework tailored to medical image annotation, for incorporating vision foundational models into the medical image annotation landscape. Hu et al. (Hu and Li, 2023) propose a SAM-based approach for multi-phase liver tumor segmentation. Their qualitative findings highlight SAM’s efficacy as a robust annotation tool within the interactive medical image segmentation community. Meanwhile, He et al. (He et al., 2023a) summarize SAM’s performance across 12 public medical image segmentation datasets, encompassing diverse organs, image modalities, and health conditions. Their findings indicate that pre-trained SAM without fine-tuning medical image data falls short of the performance of other deep learning counterparts trained. Additionally, Mazurowski et al. (Mazurowski et al., 2023) directly evaluate SAM’s capabilities on 19 distinct medical imaging datasets. This comprehensive assessment reveals SAM’s promising zero-shot medical segmentation performance.
Although SAM is a promising segmentation-focused FM, the inherent domain gap between general vision and medical imaging remains to be addressed. Huang et al. (Huang et al., 2023b) explores various evaluation strategies for SAM. Their analysis delves into the diverse factors affecting SAM’s segmentation performance, unraveling insights crucial for refining its application within the medical context. The outcome consistency and reliability of directly using SAM in medical image segmentation still need to be improved by properly refining FMs (Zhang and Jiao, 2023). Several studies have focused on adapting FMs to the complexities of medical data, aiming to improve the segmentation performance, which potentially benefits the image annotation performance. Zhang et al. (Zhang et al., 2023f) propose SAMAug to involve input augmentation, where prior maps are integrated with raw images to amplify the effectiveness of medical image segmentation. MedSAM exemplifies the process of finetuning SAM on medical images (Ma and Wang, 2023), which ultimately aims to create a versatile tool applicable across a wide range of medical tasks. Similarly, Med SAM Adapter (Wu et al., 2023b) introduces a practice approach by infusing medical-specific domain knowledge into the segmentation model. Such integration significantly enhances the model’s performance within medical image annotation tasks. Extending the range of SAM applications, Poly-SAM (Li et al., 2023b) focuses on polyp segmentation. By fine-tuning the SAM model and comparing transfer learning strategies, this study highlights the substantial potential of adapting SAM to intricate medical image segmentation tasks. Moreover, Hu et al. (Hu et al., 2023b) enhance SAM’s performance through efficient modifications by refining the lightweight task-specific prediction head while keeping the SAM encoder unaltered. Empirical validations indicate that finetuning SAM can significantly elevate its performance on medical image datasets.
Technical challenges remain in adapting SAM to the medical domain efficiently and scalably. Medical images, such as X-rays, MRIs, CT scans, and histopathological slides, could reveal domain-specific clinical patterns that require careful interpretation beyond the scope of general-purpose SAM. In particular, SAM pre-trained in general vision lacks the specialized understanding of medical anatomy, pathology, and radiology necessary for precise image analysis. Data- or parameter-efficient fine-tuning is effective while it is not scalable. Moving forward, more research efforts are required towards a more scalable medical knowledge transfer pattern for better utilizing and adapting SAM or other general-purpose FMs in medical image and text annotation.
7. Data Privacy
Safeguarding healthcare data privacy is crucial given their critical roles in clinical research, diagnosis, treatment, and disease prevention (Gkoulalas-Divanis et al., 2014). Laws and regulations have been enacted to ensure the protection of these highly sensitive personal records (Clayton et al., 2023; Malin et al., 2010), fostering the trust between patients and medical institutions.
Traditionally, healthcare data is well protected to preserve privacy by strategies such as cryptographic methods and access control mechanisms (Gertz and Jajodia, 2007; Aggarwal and Yu, 2008). Nevertheless, these approaches encounter limitations in providing absolute patient anonymity, whereby the private and confidential information of patients remains undisclosed, particularly when disseminating patient data (Fung et al., 2010). The challenge arises due to the necessity of distributing data to a broad and potentially unidentified group of recipients. To solve this, healthcare data is typically altered to remove private information before being used (de Wolf, 2012; Willenborg and De Waal, 2012; Kosorok and Laber, 2019). Perturbation methods are used to add controlled noise or randomness to data to protect privacy while allowing data analysis (Turgay et al., 2023). These methods are leveraged at a reduced level of intensity to mitigate the potential deterioration in the integrity of subsequent secondary data analyses (Reiter, 2012). However, with the emergence of accessible digital data and powerful computational resources, attackers can now obtain huge amounts of information to link released records to external data files (Reiter, 2019). This critical concern leads the research to generate synthetic data that contain no actual values of individual information (Reiter, 2023).
The advent of data-centric FMs introduces both opportunities and challenges in healthcare data privacy. On one hand, FMs offer improved capabilities for data generation, providing unidentified synthetic data (Tang et al., 2023a; Shibata et al., 2023). On the other hand, FMs tend to memorize their training data due to their massive parameters, resulting in the generation of similar outputs that potentially contain private patient information (Zhou et al., 2023b; Carlini et al., 2021).
FMs offer a practical opportunity for addressing healthcare data privacy concerns. The growing use of synthetic data represents a solution to protect private information that is in line with the key promise of FMs (Hu et al., 2023a). Deep generative models can now produce high-quality synthetic medical data with similar characteristics to the original data but without identical information (Naresh et al., 2023), thus preserving the functionality while avoiding risks of privacy leakage. For instance, recent advances in FM-based data generation methods have enabled the anonymization of medical images and texts, supporting data privacy while maintaining a wide utility (Dong et al., 2022b). FMs also enable the generation of differentially private data for model training (Sei et al., 2024), which ensures the effectiveness of model training, while the generated synthetic data preserves data privacy. Shibata et al. (Shibata et al., 2023) propose a large-scale diffusion model capable of unconditionally generating high-resolution volumetric medical images, which removes critical information while remaining trainable features, preserving privacy when sharing data in various domains beyond healthcare. Similarly, the use of synthetic imaging data has shown its promise in protecting patient records while retaining high-level segmentation performance in a multi-center evaluation (Chang et al., 2023c).
Although FMs shed light on addressing medical data privacy issues, we have recognized that FMs themselves have brought up emerging privacy protection challenges. Numerous studies have extensively investigated FMs of their propensity to memorize training data across different domains, a phenomenon referred to as memorization (Carlini et al., 2019; Thakkar et al., 2021; Zhang et al., 2021b; Carlini et al., 2022; Kandpal et al., 2022; Lee et al., 2021). FMs have been trained on private datasets (Carlini et al., 2023). This characteristic of FMs makes them prone to generating text sequences or other information directly from their training data (Carlini et al., 2021), which may include sensitive patient information. Zhou et al. (Zhou et al., 2023b) highlight the possibility of retrieving specific training samples by querying the massive FMs, leading to private information leakage. Even worse, due to the billion parameters of FMs, they are more likely to retain a greater amount of private information, making the larger model more vulnerable than smaller models. Recently, more researchers have paid attention to FM privacy issues in healthcare. The memorization phenomenon poses a risk of capturing and potentially revealing highly sensitive information, such as protected health information (Lehman et al., 2021; Pan et al., 2020). Sallam et al. (Sallam et al., 2023) highlight data privacy concerns associated with ChatGPT-generated content in healthcare education, emphasizing the need for strict regulation and full protection of patient information. To mitigate the risks of privacy breaches, it is essential to consider implementing safeguards when utilizing FMs in healthcare. These safeguards should encompass various stages such as data processing, model training, model inference, and system deployment.
To meet the growing need for data privacy protection across diverse contexts, FM-enabled data protection and privacy preservation requires thorough research investigation. (1) To overcome the potential patient information leakage in the model training, federated learning naturally provides a distributed network solution for protecting in-house patient data in local institutions. This training strategy could ensure FM is trained locally without releasing the in-house patient data for FM training and updating. (2) The design of FM architectures is expected to improve the capability of detecting risk questions that may leak patient information instead of providing answers without considering any privacy concerns. Once the FMs are able to detect the questions that may expose patient identifications, the defense mechanism in FM architecture design should perform as patient information protectors to anonymize the unique characteristics of patients. Research efforts will continue to focus on FM architecture design and deployment for preventing patient identification from being recovered by input queries, while protecting sensitive personal information leakage.
8. Performance Evaluation
Systematic and reliable evaluation is critical for assessing the performance and safety of AI models deployed in healthcare settings. The evaluation of FMs is challenging owing to their extensive utilization given their own model scale and complexity (Chang et al., 2023b). In this section, we discuss FM evaluation strategies and challenges from three key aspects, including benchmarking (Section 8.1), human evaluation (Section 8.2), and automated evaluation (Section 8.3).
Figure 6. Foundation model evaluation strategies.
8.1. Benchmarking
Researchers have established benchmarks for model evaluation in the healthcare field, inspired by the progress of AI research in vision and language domains (Mincu and Roy, 2022; Johnson et al., 2016; Gu et al., 2021). A remarkable feature is that these benchmarks often build on sample diversity and large quantity scale to facilitate FM evaluation. For example, The Biomedical Language Understanding Evaluation benchmark (BLUE) (Peng et al., 2019) covers five language tasks with ten biomedical and clinical text datasets, mixing PubMed and MIMIC-III (Johnson et al., 2016) based applications. The Biomedical Language Understanding & Reasoning Benchmark (BLURB) (Gu et al., 2021) focuses on biomedical datasets and improves BLUE by including question-answering tasks which are frequently employed to evaluate large language models (LLMs). Table 1 displays a representative list of public medical databases, benchmarks, and datasets that can facilitate FM training, inference, and evaluation.
Table 1. Public medical databases, benchmarks, and datasets that can facilitate foundation model development and application in medicine.
Database/Benchmark/Dataset Description Task type MIMIC-III (Johnson et al., 2016) Critical care data for over 40,000 patients Text BLUE (Peng et al., 2019) 5 language tasks with 10 biomedical and clinical text datasets Text MedMentions (Mohan and Li, 2019) 4,392 papers annotated by experts with mentions of UMLS entities Text webMedQA (He et al., 2019) 63,284 real-world Chinese medical questions with 300K answers Text PubMedQA (Jin et al., 2019) 1K expert-annotated, 61.2K unlabeled, and 211.3K artificially generated biomedical QA instances Text BLURB (Gu et al., 2021) 13 biomedical NLP datasets in 6 tasks Text CBLUE (Zhang et al., 2021a) A Chinese medical NLP benchmark with 18 datasets Text MedQA-USMLE (Jin et al., 2021) 61,097 multiple choice questions based on USMLE in three languages Text MedMCQA (Pal et al., 2022) 194K multiple-choice questions covering 2.4K healthcare topics Text MultiMedQA (Singhal et al., 2022) 6 existing and 1 online-collected medical QA dataset Text ClinicalTrials.gov [^3] https://clinicaltrials.gov/ An online database of clinical research studies, including clinical trials and observational studies Text Medical Meadow (Han et al., 2023) 16M medical QA pairs collected from 9 sources Text Huatuo-26M (Li et al., 2023d) 26M Chinese medical QA pairs Text ChestXray-NIHCC (Wang et al., 2017) 100K radiographs with labels from more than 30,000 patients Imaging DeepLesion (Yan et al., 2018) 32K CT scans with annotations and semantic labels from radiological reports Imaging Kather Colon Dataset (Kather et al., 2018) 100K histological images of human colorectal cancer and healthy tissue Imaging CheXpert (Irvin et al., 2019) 224,316 chest radiographs of 65,240 patients Imaging EchoNet-Dynamic (Ouyang et al., 2020) 10,030 expert-annotated echocardiogram videos Imaging ISIC [^4] https://www.isic-archive.com/ An archive containing 23K skin lesion images with labels Imaging Med-MNIST v2 (Yang et al., 2023a) 12 2D and 6 3D datasets for biomedical image classification Imaging AbdomenAtlas-8K (Qu et al., 2023) 8,448 CT volumes with per-voxel annotated eight abdominal organs Imaging Virchow (Vorontsov et al., 2023) 1.5M pathological scans from 120K patients Imaging RETFound (Zhou et al., 2023a) Unannotated retinal images, containing 904,170 CFPs and 736,442 OCT scans Imaging dbSNP (Sherry et al., 2001) A collection of human single nucleotide variations, microsatellites, and small-scale insertions and deletions Genomics ENCODE (Consortium et al., 2012) A platform of genomics data and encyclopedia with integrative-level and ground-level annotations Genomics 1000 Genomes Project (Consortium et al., 2015) A comprehensive catalog of human genetic variations Genomics ChEMBL (Gaulton et al., 2012) 20M bioactivity measurements for 2.4M distinct compounds and 15K protein targets Drug DrugBank (Wishart et al., 2018) A web-enabled structured database of molecular information about drugs Drug PubChem (Kim et al., 2023) A collection of 900+ sources of chemical information data Drug TCGA [^5] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A landmark cancer genomics program, molecularly characterized over 20,000 primary cancer and matched normal samples spanning 33 cancer types Multi-modal MIMIC-CXR (Johnson et al., 2019) 227,835 chest imaging studies with free-text reports for 65,379 patients Multi-modal MIMIC-IV (Johnson et al., 2020) Clinical information for hospital stays of over 60,000 patients Multi-modal SwissProtCLAP (Liu et al., 2023g) 441K text-protein sequence pairs Multi-modal PMC-VQA (Zhang et al., 2023d) 227K VQA pairs of 149K images of various modalities or diseases Multi-modal MedMD (Wu et al., 2023e) 15.5M 2D scans and 180k 3D radiology scans with textual descriptions Multi-modal PathCap (Sun et al., 2023) 142K pathology image-caption pairs from various sources Multi-modal PathInstruct (Sun et al., 2023) 180K samples of LLM-generated instruction-following data Multi-modal OpenPath (Huang et al., 2023a) 208,414 pathology images paired with natural language descriptions Multi-modal Quilt-1M (Ikezogwo et al., 2023) 1M image-text pairs for histopathology Multi-modal PMC-OA (Lin et al., 2023b) 1.6M fine-grained biomedical image-text pairs Multi-modal
Current benchmarks are largely focused on static datasets, which are unable to capture the full complexities of how FMs behave in dynamic environments with continuous variants, real-time adjustments, and interaction with humans. In addition, FMs are more likely to have seen the common benchmarks during pre-training, which can potentially lead to biased performance during evaluation (Shi et al., 2023). From a data-centric perspective, developing dynamic or temporal datasets that reflect complex real-world scenarios could provide a more accurate representation of FM behavior in healthcare applications.
Benchmarks also cover different evaluation metrics for a wide range of tasks. Conventional evaluation metrics include accuracy, area under the curve (AUC), and mean average precision (mAP) for classification, intersection over union (IoU) for image segmentation, and bilingual evaluation understudy (BLEU) and BERTScore (Zhang et al., 2019) for text-related tasks, which mostly emphasize precise model predictions. However, FM evaluation metrics largely go beyond mere accuracy. Recent work has queried whether benchmark evaluation metrics align with human values (Wornow et al., 2023; He et al., 2023b). The lack of such alignment could pose misleading estimates of FM capabilities leading to harmful results. In fact, LLM alignment evaluation from perspectives of reliability, safety, fairness, explainability, and robustness underscores the importance of LLM trustworthiness (Liu et al., 2023d). Yet we lack substantial research benchmarks for FM trustworthiness evaluation in the healthcare field.
8.2. Human Evaluation
In real-world healthcare applications, there are often scenarios that are not covered by the major benchmarks (e.g., rare disease assessment). Thus, a critical approach to FM evaluation is to consult human experts. For example, Chambon et al. (Chambon et al., 2022) assess the clinical correctness of FM-generated chest X-ray images with the help of radiologists. Lyu et al. (Lyu et al., 2023) invite two experienced radiologists to evaluate the overall score, completeness, and correctness of LLM-translated radiology reports. Peng et al. (Peng et al., 2023) invite two physician evaluators to perform Turing evaluation of 30 paragraphs written by FMs and humans respectively to assess the readability and clinical relevancy of FM-generated clinical text. Singhal et al. (Singhal et al., 2022) present a framework for human evaluation of LLM answers to medical questions. Moor et al. (Moor et al., 2023b) implement a human evaluation application for clinical experts to rate the LLM-generated answers for medical image-related questions.
Introducing human evaluation to the data curation and model training process can make model outcomes more aligned with human values. For instance, the methodology of Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) combines human evaluation with reinforcement learning techniques to train language models that exhibit strong alignment with user instructions, which serves as the cornerstone of modern chatbots like ChatGPT (OpenAI, 2021). The authors also adopt human ranking for evaluating the final model performance. Similarly in the healthcare field, IvyGPT (Wang et al., 2023b) provides rich diagnosis and treatment answers by applying the RLHF technique on medical question-answering data. However, relying on human experts’ assessment can incur significant costs and introduce subjectivity and bias, particularly given the flexibility of FM outputs. Training and hiring healthcare experts can be expensive (Van Smeden et al., 2022). Moreover, variations in experts’ backgrounds, prior experience, motivations, and methods of inquiry make uniform assessment difficult to achieve in practice (Gehrmann et al., 2023).
8.3. Automated Evaluation
Leveraging powerful FMs for automated evaluation is a promising research direction. A high-performing and well-aligned FM can be employed as a benchmark or an evaluation tool to automatically assess other FMs, eliminating the need of static benchmarks or human experts. For instance, Chen et al. (Chen et al., 2023c) explore reference-free evaluation methods that prompt ChatGPT to score the quality of model-generated texts without a pre-defined ground truth. Chiang et al. (Chiang and Lee, 2023) show that LLM evaluation can be stable over prompt formatting and consistent with human experts. Jain et al. (Jain et al., 2023) propose self-supervised evaluation strategies to assess LLM properties without benchmarks or human annotations. Ye et al. propose FLASK (Ye et al., 2023), a fine-grained LLM evaluation protocol that decomposes coarse-level scoring to instance-wise skill assessment. FLASK prompts GPT-4 to assign a score with a rationale for each alignment skill of the target FM-generated text instance. From a data-centric perspective, the benchmark FM with extensive knowledge and human alignment embedded in its massive pre-training data can evaluate the performance of other FMs.
In the realm of healthcare, leveraging FMs for evaluation has not yet been fully explored. FMs have demonstrated the ability to understand and generate medical text or images (Nori et al., 2023; Roy et al., 2023; Yu et al., 2023). Nevertheless, they have shown limitations including inconsistent performance across datasets (Cheng et al., 2023b) and lower accuracy compared with state-of-the-art in-domain learning methods (He et al., 2023a). Whether automated evaluation via FMs is trustworthy for healthcare applications remains an unresolved question. An integrated approach that combines automated evaluation with human expert supervision could potentially yield more reliable results. Together, we expect the development of more evaluation benchmarks (for imaging, texts, or genomics) within the healthcare field with rigorous design, deployment, and validation.
9. Challenges and Opportunities
Despite the promise of FMs in healthcare, open challenges remain in developing and adapting trustworthy FMs to gain insights into patient outcome and clinical workflow. We discuss key directions for building reliable healthcare-focused FMs towards better human-AI alignments, addressing various aspects in hallucination, bias, and proper regulation protocols.
9.1. Healthcare-Focused Foundation Model Development
Current general-purpose FMs must address myriad challenges for extracting reliable insights given the domain gap between general and healthcare data. Different from natural images and texts, medical images (e.g., pathological scans) and semantic text profiles (e.g., clinical reports) can reveal unique disease patterns of patients. Training an FM from scratch for a particular medical task is computationally expensive for scientific research laboratories. As a result, fine-tuning and adapters are increasingly considered (Zhang and Metaxas, 2023). These data- or parameter-efficient tuning strategies provide means for extracting domain-specific representations using FMs. Yet, such tuning strategies often come at the risk of catastrophic forgetting (Ramasesh et al., 2021), which refers to the model forgetting the knowledge obtained from pre-training data during the process of fine-tuning. Therefore, benchmarking these fine-tuning and prompt engineering strategies in the healthcare domain is much required to guide the fair use of FMs in various healthcare settings.
Healthcare information integration via FM becomes a prioritized challenge for assisting clinical routines. We expect a remarkable growth in adapting FMs to capture multi-scale healthcare insights. Multi-modal data synergy could greatly capture multi-scale patterns toward better disease understanding, and therefore developing multi-scale FM architectures is highly considerable (Gao et al., 2023). For example, genomics sequencing profiles could reveal complex molecular patterns, which are valuable for determining targeted therapy and disease classification. Meanwhile, histopathological images contain comprehensive tissue microenvironment information, enabling precise disease diagnosis and prognosis. Integrating these biological data could open perspectives for exploring the key insights of genotype-phenotype interactions across modalities for better patient outcome prediction and management (Ding et al., 2023b).
9.2. Hallucination
Hallucination in the context of FM refers to a circumstance where large language models (LLMs) generate plausible yet inaccurate content (Umapathi et al., 2023). The generated content (e.g., text details, facts, or claims) is fictional, misleading, or even fabricated rather than providing truthful information. Hallucination can be caused by several factors in the training data, including issues related to quality, scale, and internal bias. Mitigating hallucination in healthcare FM applications is important because misinformation can cause severe consequences for all healthcare stakeholders. For instance, the hallucination-affected FMs could produce content that can potentially affect healthcare diagnosis, decision-making, and patient care. To address this challenge, a critical step is to properly evaluate the severity of the hallucination. The evaluation metrics for detecting hallucination should consider key factors such as factual accuracy, coherence, and consistency (Zhang et al., 2023b). The benchmark on Medical Domain Hallucination Test (Med-HALT) (Umapathi et al., 2023) is an example for evaluating hallucination in LLMs. Med-HALT includes reasoning and memory-based hallucination tests for assessing the problem-solving and information-retrieval capabilities of LLMs in medical contexts. Furthermore, human-AI collaboration could be promising for reducing hallucination in FMs. Crowdsourcing platforms can also be used to gather human assessments of FM-generated content (Rawte et al., 2023). Finally, adversarial testing for healthcare-related FMs is helpful in identifying the hallucinated contents to improve FMs’ trustworthiness and generalization (Rawte et al., 2023; Zhang et al., 2023b).
9.3. Bias
Bias represents misrepresentation, attribution error, or factual distortions in the cycle of FM development, resulting in inequities, stereotyping, and discriminatory consequences (Ferrara, 2023). Mitigating these biases could reduce misunderstanding of the generated knowledge and enhance the trustworthiness of FM. For example, based on the performance of evaluating ChatGPT on the United States Medical Licensing Examination (USMLE), studies (Kung et al., 2023; Wang et al., 2023d) found that ChatGPT could have a better performance in English than Asian-linguistic contexts. A major reason is that linguistic bias remains in the LLM training process. Thus addressing the cultural, linguistic, demographic, and political biases in FM training data becomes a priority (Ferrara, 2023). In addition, the rigorous audit scheme in the process of dataset curation could alleviate the harmful information that may already exist in the massive training data.
From a model perspective, the inherent bias in FM architectures should be recognized by healthcare stakeholders and developers. Currently, FM architectures and training patterns lack the defense and detection capability against adversarial manipulation. FMs could propagate all knowledge learned from training data without any regulation of biased knowledge detection. This vulnerability can potentially hurt healthcare stakeholders by generating false or harmful information. Meanwhile, adversarial attack training is promising to enhance the harmful information defense and detection capability (Ferrara, 2023). Furthermore, the bias of human-AI alignment remains in FM development, where matching human value and model output will be a long-term challenge. To mitigate this bias, strong guidance from human experts in dataset curation and FM outcome evaluation is necessary in the healthcare domain. Human experts can provide reliable feedback to FM developers for identifying and addressing bias-related issues (Ferrara, 2023).
9.4. Regulation
AI model governance and real-world utility are essential for healthcare stakeholders to deploy, assess, and apply FMs in different settings. Yet current efforts have not complied with HIPAA or governmental regulations (e.g., FDA) to provide on-demand clinical service. With the increasing use of healthcare AI, FDA started regulating software as a medical device that software solutions performing medical functions in the prevention, diagnosis, treatment, or monitoring of various diseases (Gilbert et al., 2023; Meskó and Topol, 2023). We expect that FM efforts will be viewed as a new form of medical device for healthcare stakeholders. To highlight, the emerging workflow of FM-enabled medical devices could be considered in the following regulatory directions (Gilbert et al., 2023): (1) We need to define the purpose and utility scope of FMs in real-world clinical applications. (2) The performance of FMs should be rigorously benchmarked on the authoritative clinical data sources. (3) The user-accessible guideline should be documented for safe and trustworthy usage. (4) FMs should be applied in comprehensive clinical trials to demonstrate their efficacy. Finally, continuous regulation is required even after FM deployment as we often deal with changing tasks and settings (Moor et al., 2023a).
The open-source community is highly valuable for FM development, revealing insights into training data, model architecture, hyperparameter settings, and user interactions. These fine details could help establish strong regulation mechanisms to enhance the trustworthy use of FMs in healthcare. To do so, medical data standardization is essential during the dataset preparation. The standardization will ensure the data collected from multiple institutions are organized in a structured format compatible with FM model development to avoid data inconsistencies. For instance, the Digital Imaging and Communications in Medicine (DICOM) standards and the Picture Archiving and Communication System (PACS) provide a standardized platform for imaging data management. Also, the CEDAR Workbench (Musen et al., 2022) is a web-based platform allowing us to manage and share the library of metadata templates. To enhance the sharing of scientific data, the FAIR (Findable, Accessible, Interoperable, and Reproducible) principle (Wilkinson et al., 2016) provides consistent metadata preparation. Popularizing the standardized guideline documents, such as the ”datasheet” (Gebru et al., 2021), is crucial to simplify metadata preparation. These efforts are helpful for us to understand the comprehensive details of the dataset, enhancing the wide adoption of clinical data to accelerate the progress of FM research in healthcare.
10. Conclusions
The striking progress of foundation models (FMs) and their applications in healthcare open up possibilities for better patient management and efficient clinical workflow. In these efforts, collecting, processing, and analyzing scalable medical data has become increasingly crucial for FM research. In this survey, we have offered an overview of FM challenges from a data-centric perspective. FMs possess great potential to mitigate data challenges in healthcare, including data imbalance and bias, data scarcity, and high annotation costs. Due to FM’s strong content generation capabilities, there is a remarkable need for greater vigilance regarding data privacy, data bias, and ethical considerations about the generated medical knowledge. Only by adequately and reliably addressing the data-centric challenges can we better leverage the power of FMs across a broader scope of medicine and healthcare.
Appendix A Healthcare Data Modalities
There are a plethora of healthcare data modalities collected in different clinical settings, such as imaging, clinical notes, biosensor records, and electroencephalography. These data records provide multi-scale information for professionals in clinical diagnosis, prognosis, and treatment development (Azam et al., 2022; Wang and Summers, 2012; Kumar et al., 2020, 2019; Subramanian et al., 2020; Heikenfeld et al., 2018). Patient diagnosis is routinely based on analyzing complex disease patterns derived from various healthcare data modalities. The diversity of medical modalities comes from the distinct methods of data acquisition, including the use of invasive procedures or the types of used medical devices. We discuss common healthcare data modalities in clinical practices, including radiology, histopathology, molecular information, clinical information, blood testing, biosensors, and electroencephalography.
Radiology commonly includes X-ray, magnetic resonance imaging (MRI), computed tomography (CT), positron emission tomography (PET), and ultrasound scans. From radiological image inputs, we are able to capture visual patterns that are helpful for clinicians to make diagnoses, providing valuable information about the anatomical structure and function of the body’s internal organs and tissues (Azam et al., 2022; Wang and Summers, 2012). Various imaging principles, such as ionizing radiation, radio waves, gamma rays, and sound waves, are utilized to visualize the internal structures and activities non-invasively (Herring, 2019). These radiographs provide key information about the body’s tissues, organs, and bones as well as metabolic processes that are crucial to assess disease characteristics. Radiological images play an important role in the application of FMs in healthcare (Sowrirajan et al., 2021; Tiu et al., 2022). In particular, radiological images and reports can be analyzed in a joint-modal manner, providing a more comprehensive understanding of patient conditions (Chambon et al., 2022; Zhang et al., 2022).
Histopathology involves examining visually perceptible changes in cells and tissues to study the manifestations of diseases (Kumar et al., 2020, 2019). In clinical practice, tissue is first removed from the body, and then the histological sections are placed onto glass slides. A high-resolution slide scanner equipped with microscope optics serially captures image tiles across the tissue section. The seamless reconstruction of these tile images yields a single monolithic digital representation of the entire glass slide, typically within the order of billions of pixels. Staining techniques highlight architectural and morphological details, demonstrating the changes in cells and the actual causes of the illness. Therefore, the microscopic examination of stained tissue sections offers a definitive disease diagnosis. Different magnifications in histopathology imaging can reveal morphological details, offering multi-level visual perceptions in clinical scenarios. Higher magnifications reveal finer attributes, while lower magnifications facilitate the understanding of overall tissue morphology (Ashtaiwi, 2022). Due to the high resolution of histopathological images, current FMs mostly take split image patches as input instead of whole slides (Huang et al., 2023a; Lu et al., 2023a).
Molecular information typically encompasses various categories of biomarkers, including DNA, RNA, proteins, metabolites, and more recently microbiome profiles (Mishra and Verma, 2010). These molecular signatures have remarkably aided disease diagnosis, prognosis, treatment, and monitoring of therapeutic response across cancer, infectious disease, and genetic disorders. Molecular diagnostics analyzes biofluids or tissue samples to identify abnormalities at the genetic and molecular level, which allow disease signature detection and target therapy planning (Subramanian et al., 2020; Lu et al., 2005). However, interpreting complex multidimensional molecular data requires specialized bioinformatics pipelines, particularly from the perspective of FM (Cui et al., 2023b). While molecular profiles offer valuable insights into disease formation and progression, data-driven analytics and clinical trial validation are increasingly necessary before the widespread deployment of molecular testing into real-world healthcare utility.
Clinical information typically encompasses structured metadata documenting patients’ demographic details, medical history, interventions, outcomes, and other relevant variables (Simpao et al., 2014). Sources of clinical data include electronic health records (EHR), clinical trial case report forms, disease registries, and health surveys. These structured fields capture details such as age, gender, medications, procedures, family history, social factors, and clinician notes. Such metadata complements diagnostic results to provide crucial contexts for enhancing patient care. Clinical information can aid patient stratification, disease pattern recognition, healthcare predictive analytics, and decision-making optimization in precision medicine applications (Shaikhina et al., 2019). Meanwhile, clinical text and imaging records stand as the frontier data modalities for FM development. This underscores the importance of clinical information in the advancement of healthcare FMs (Alsentzer et al., 2019). However, heterogeneity across institutions and populations poses data integration challenges, such as the documentation conventions of various institutions as well as individual writing styles. Standardization of clinical terminologies and ontologies is thus essential for unambiguous data sharing and mining (Kim et al., 2019). As a result, multi-scale clinical data records could complement other data modalities to enable a holistic understanding of patients and personalized therapies.
Blood testing reveals rich information by analyzing constituents of blood specimens like cell counts, proteins, metabolites, and hormones in patient blood samples (Burtis and Bruns, 2014). The specimens are usually obtained through minimally invasive venipuncture or fingerprick and are analyzed using automated systems. Analysis techniques encompass microscopy, flow cytometry, chromatography, etc (Li HY, 2022). The analysis of blood testing for patients is usually based on the assessment of biomarker level, which reflects patient health status and assists in screening, diagnosis, treatment, and monitoring across a wide range of conditions (Brusselle et al., 2018; Hier et al., 2021). AI techniques have been explored for the discovery of blood biomarkers in cancer (Visaggi et al., 2021). Another example is MediTab (Wang et al., 2023c), an FM for medical tabular data prediction that scales up model pre-training with blood testing data. However, FM for tabular data analytics including blood testing data remains challenging due to the difference between the structured data and free-form text (Bisercic et al., 2023).
Biosensors are devices or probes containing a biological recognition element coupled to a transducer that converts a physiological or biochemical signal into a measurable electronic output (Haleem et al., 2021). Biosensing techniques typically include electrochemical, optical, piezoelectric, thermal, and magnetic detection (Naresh and Lee, 2021). Biosensors allow continuous tracking of vital signs, physiological signals, and biomarkers by transducing biological responses into electrical outputs (Heikenfeld et al., 2018). These multi-modal data streams from devices monitoring heart rates and glucose levels could enable personalized medicine through remote patient surveillance and assessment of health status changes (Majumder et al., 2017). In particular, biosensors offer great promise for ubiquitous real-time health monitoring, thus providing extra information in multi-modal computer-aided diagnosis.
Electroencephalography (EEG) measures electrical brain activity by recording voltage fluctuations resulting from ionic currents within neurons (Steriade, 2005). Electrodes are placed on the scalp to record electrical activity emerging from firing neurons in the brain that manifest as brain waves with distinct rhythmic patterns (Van Diessen et al., 2015). Analysis of brain waves can facilitate the diagnosis of neurological disorders, concussions, sleep abnormalities, and other conditions (Craik et al., 2019). Wearable EEG devices can facilitate longitudinal brain monitoring to assess disease progression, recovery, and treatment effects. Research has demonstrated that self-supervised EEG representation learning with massive unlabeled EEG signals for few-shot sleep staging tasks is feasible (Yang et al., 2023b). However, signal artifacts, low spatial resolution, and inter-subject variability are ongoing research challenges. Overall, EEG signals offer a non-invasive access to correlate brain physiology with function, behavior, and pathology, which greatly facilitates clinical diagnosis and healthcare research.
Appendix B Healthcare Foundation Models
Table 2 highlights healthcare FMs, including their model structure, initialization, pre-training data, and the link reference to the project.
Table 2. Typical foundation models in healthcare and medicine.
Model Base Initial Model Weights Pre-training Data SciBERT (Beltagy et al., 2019) BERT (Devlin et al., 2018) - Semantic Scholar (Ammar et al., 2018) Clinical BERT (Alsentzer et al., 2019) BERT BERT MIMIC-III (Johnson et al., 2016) BlueBERT (Peng et al., 2019) BERT - PubMed [^6] https://pubmed.ncbi.nlm.nih.gov/ + MIMIC-III BioBERT (Lee et al., 2020) BERT BERT PubMed + PMC [^7] https://www.ncbi.nlm.nih.gov/pmc/ PubMedBERT (Gu et al., 2021) BERT - PubMed BioLinkBERT (Yasunaga et al., 2022) BERT - PubMed BioGPT (Luo et al., 2022) GPT-2 (Radford et al., 2019) - PubMed Clinical-T5 (Lehman and Johnson, 2023) T5 (Raffel et al., 2020) T5 MIMIC-III + MIMIC-IV (Johnson et al., 2020) HIPT (Chen et al., 2022) DINO (Caron et al., 2021) - TCGA [^8] https://www.cancer.gov/ccg/research/genome-sequencing/tcga CTransPath (Wang et al., 2022d) SRCL (Wang et al., 2022d) - TCGA + PAIP [^9] http://www.wisepaip.org/paip/ RETFound (Zhou et al., 2023a) MAE (He et al., 2022) Vision Transformer (Dosovitskiy et al., 2020) RETFound REMEDIS (Azizi et al., 2023) SimCLR (Chen et al., 2020) BiT (Beyer et al., 2022) MIMIC-IV + CheXpert (Irvin et al., 2019) Virchow (Vorontsov et al., 2023) DINOv2 (Oquab et al., 2023) - Virchow PubMedCLIP (Eslami et al., 2021) CLIP (Radford et al., 2021) CLIP ROCO (Pelka et al., 2018) CheXzero (Tiu et al., 2022) CLIP CLIP MIMIC-CXR (Johnson et al., 2019) MedCLIP (Wang et al., 2022c) CLIP Clinical BERT + SwinTransformer (Liu et al., 2021) CheXpert + MIMIC-CXR BiomedCLIP (Zhang et al., 2023e) CLIP PubMedBERT PMC-15M (Zhang et al., 2023e) PMC-CLIP (Lin et al., 2023b) CLIP PubMedBERT PMC-OA (Lin et al., 2023b) MI-Zero (Lu et al., 2023b) CLIP HistPathGPT (Lu et al., 2023b) + CTransPath (Wang et al., 2022e) ARCH (Gamper and Rajpoot, 2021) PLIP (Huang et al., 2023a) CLIP CLIP OpenPath (Huang et al., 2023a) QuiltNet (Ikezogwo et al., 2023) CLIP CLIP Quilt-1M (Ikezogwo et al., 2023) KAD (Zhang et al., 2023c) CLIP - MIMIC-CXR + UMLS (Bodenreider, 2004) CONCH (Lu et al., 2023a) CoCa (Yu et al., 2022) - PubMed + PMC Med-Flamingo (Moor et al., 2023b) Flamingo (Alayrac et al., 2022) Flamingo MTB (Moor et al., 2023b) + PMC-OA Med-PaLM (Singhal et al., 2022) PaLM (Chowdhery et al., 2022) PaLM MultiMedQA (Singhal et al., 2022) ChatDoctor (Yunxiang et al., 2023) LLaMA (Touvron et al., 2023a) LLaMA HealthCareMagic-100k (Yunxiang et al., 2023) PMC-LLaMA (Wu et al., 2023d) LLaMA LLaMA MedC (Wu et al., 2023d) Clinical LLaMA-LoRA (Gema et al., 2023) LLaMA LLaMA / PMC-LLaMA MIMIC-IV (Johnson et al., 2020) BioMedGPT (Luo et al., 2023) LLaMA 2 (Touvron et al., 2023b) LLaMA 2 S2ORC (Lo et al., 2019) LLaVA-Med (Li et al., 2023e) LLaVA (Liu et al., 2023b) LLaVA PMC-15M (Zhang et al., 2023e) + GPT-4 (OpenAI, 2023) Qilin-Med-VL (Liu et al., 2023c) LLaVA Chinese-LLaMA2 + CLIP ChiMed-VL (Liu et al., 2023c) MedVInT (Zhang et al., 2023d) - PMC-LLaMA + PMC-CLIP PMC-VQA (Zhang et al., 2023d) RadFM (Wu et al., 2023e) - PMC-LLaMA MedMD (Wu et al., 2023e) ESM-2 (Lin et al., 2023a) Transformer (Vaswani et al., 2017) - UniRef (Suzek et al., 2015) GET (Fu et al., 2023) Transformer - GET AlphaMissense (Cheng et al., 2023a) AlphaFold (Jumper et al., 2021) - PDB (Berman et al., 2000) + UniRef

### v2

@@
 In the healthcare system, collecting high-quality records could enable a comprehensive understanding of patient characteristics (imaging, genomics, and lab testing data) (Johnson and Khoshgoftaar, 2023; Singh, 2023; Alfasly et al., 2023).
-As illustrated, data-centric strategies promise to reshape clinical workflow (Rao et al., 2023; Juluru et al., 2021), enable precise diagnosis (Hunter et al., 2022), and uncover insights into treatment (Chen et al., 2023a).
+As illustrated, data-centric strategies promise to reshape clinical workflow (Rao et al., 2023; Juluru et al., 2021), enable precise diagnosis (Hunter et al., 2022), and uncover insights into treatment (Chen et al., 2023c).
 Figure 1. Data-centric foundation models in computational healthcare.
@@
 To build responsible solutions for healthcare AI, the evolving perspective on AI-human alignment (Gabriel, 2020; Ngo et al., 2022) has become increasingly important.
-We discuss the necessity of the real-world applications of FMs aligned with human ethics, equity, and societal norms to reduce potential risks in performance assessment, ethical compliance, and patient safety (Liu et al., 2023d; Ozmen Garibay et al., 2023; Liang et al., 2022; Hathaliya and Tanwar, 2020).
+We discuss the necessity of the real-world applications of FMs aligned with human ethics, equity, and societal norms to reduce potential risks in performance assessment, ethical compliance, and patient safety (Liu et al., 2023e; Ozmen Garibay et al., 2023; Liang et al., 2022; Hathaliya and Tanwar, 2020).
 In the FM era, enabling AI-human alignment further underscores the significance of data focus, motivating us to prioritize the data-centric challenges in the landscape of computational healthcare.
@@
 Pre-Training [General FMs: CLIP (Radford et al., 2021); ALIGN (Jia et al., 2021); Florence (Yuan et al., 2021);
-GLIP (Li et al., 2022); CoCa (Yu et al., 2022); FLAVA (Singh et al., 2022); CLAP (Elizalde et al., 2023)] [Medical FMs: ConVIRT (Zhang et al., 2022); BioViL (Boecking et al., 2022);
+GLIP (Li et al., 2022); CoCa (Yu et al., 2022); FLAVA (Singh et al., 2022); CLAP (Elizalde et al., 2023)] [Medical FMs: ConVIRT (Zhang et al., 2022); EchoCLIP (Christensen et al., 2024);
 PLIP (Huang et al., 2023a); PubMedCLIP (Eslami et al., 2021); CheXzero (Tiu et al., 2022);
-BiomedJourney (Gu et al., 2023)] ] [Leveraging LLMs [General FMs: Frozen (Tsimpoukelli et al., 2021); SimVLM (Wang et al., 2021);
-Flamingo (Alayrac et al., 2022); BLIP-2 (Li et al., 2023c); LLaVA (Liu et al., 2023b);
+BiomedJourney (Gu et al., 2023); BioViL (Boecking et al., 2022); PRISM (Shaikovski et al., 2024);
+MoleculeSTM (Liu et al., 2023c)] ] [Leveraging LLMs [General FMs: BLIP-2 (Li et al., 2023c); SimVLM (Wang et al., 2021);
+Flamingo (Alayrac et al., 2022); GPT-4o (OpenAI, 2024); LLaVA (Liu et al., 2023b);
 NExT-GPT (Wu et al., 2023a)] [Medical FMs: LLaVA-Med (Li et al., 2023e); MedVInT (Zhang et al., 2023d);
-Visual Med-Alpaca (Shu et al., 2023); Med-Flamingo (Moor et al., 2023b)] ] ] [Quantity
+Visual Med-Alpaca (Shu et al., 2023); Med-Flamingo (Moor et al., 2023b);
+CheXagent (Chen et al., 2024)] ] ] [Quantity
 (Section 5) [Data Augmentation [Text: ChatGPT (Tang et al., 2023a)] [Imaging: Latent Diffusion (Packhäuser et al., 2022; Pinaya et al., 2022; Chambon et al., 2022);
@@
 […]
-CLIP (Zhang et al., 2023a); GLIP (Yi et al., 2023)] [Medical FM-based: REMEDIS (Azizi et al., 2023);
+CLIP (Zhang et al., 2023a); GLIP (Yi et al., 2023); REALM (Zhu et al., 2024); BioRAG (Wang et al., 2024)] [Medical FM-based: REMEDIS (Azizi et al., 2023);
 Biomedical LMs (Mishra et al., 2023; Zhang et al., 2023a)] ] [Internet-Scale
@@
 PMC-OA (Lin et al., 2023b); HealthCareMagic-100k (Yunxiang et al., 2023)] ] ] [Annotation
-(Section 6) [Text Annotation [Gilson et al. (Gilson et al., 2022); PathAsst (Sun et al., 2023); DeID-GPT (Liu et al., 2023e);
-Med-PaLM 2 (Singhal et al., 2023); Tang et al. (Tang et al., 2023b)] ] [Image Annotation [Zero-shot SAM: SAM (Kirillov et al., 2023); Hu et al. (Hu and Li, 2023);
+(Section 6) [Text Annotation [Gilson et al. (Gilson et al., 2022); PathAsst (Sun et al., 2023); DeID-GPT (Liu et al., 2023f);
+Med-PaLM 2 (Singhal et al., 2023a); MedGraphRAG (Wu et al., 2024);
+Tang et al. (Tang et al., 2023b)] ] [Image Annotation [Zero-shot SAM: SAM (Kirillov et al., 2023); Hu et al. (Hu and Li, 2023);
 He et al. (He et al., 2023a); Mazurowski et al. (Mazurowski et al., 2023);
@@
 With the information-rich pre-training, FMs exhibit a comprehensive understanding of concepts and their interrelationships, facilitating the application of knowledge to downstream tasks through adaptation.
-As a concrete example of FM, large language models (LLMs) are pre-trained on Internet-scale corpora showing impressive semantic understanding and high-quality text generation (Brown et al., 2020; Touvron et al., 2023a; Chowdhery et al., 2022).
+As a concrete example of FM, large language models (LLMs) are pre-trained on Internet-scale corpora, showing impressive semantic understanding and high-quality text generation (Brown et al., 2020; Touvron et al., 2023a; Bai et al., 2023; OpenAI, 2023).
 In particular, LLMs emphasize the engineering design of input context, also known as the prompt.
-This design opens the door for human-and-machine interaction, allowing us to feed various prompt queries (e.g., text sequences) into a well-trained FM towards generating desired content or outcomes.
+This design opens the door for human-machine interaction, allowing us to feed various prompt queries into a well-trained FM towards generating desired content or outcomes.
 In this section, we discuss the essential concepts and useful capabilities in the cycle of general FM development, including large-scale pre-training (Section 2.1), fine-tuning (Section 2.2), and in-context learning (Section 2.3).
@@
 In detail, fine-tuning refers to updating pre-trained model parameters through forward and backward passes and gradient descent on task-specific supervised data.
-Fine-tuning is a critical technique in constructing medical domain-specific FMs (Lee et al., 2020; Gema et al., 2023; Li et al., 2023e).
+Fine-tuning is a critical technique in constructing medical domain-specific FMs (Gema et al., 2023; Li et al., 2023e; Saab et al., 2024).
 However, conventional fine-tuning poses significant challenges including high cost, extensive data requirements, and limited generalization to unseen tasks.
@@
 Low-Rank Adaptation (LoRA) (Hu et al., 2021) reparameterizes the weight matrices with multiplications of low-rank matrices.
-PEFT methods demonstrate potential effectiveness in healthcare (Gema et al., 2023).
+PEFT methods have demonstrated potential effectiveness in healthcare (Gema et al., 2023).
 2.2.2. Instruction tuning
@@
 We introduce essential ICL strategies related to FM applications.
-Zero-shot is the simplest type of prompt, where we directly assign a task to the model, e.g., “Tell me a joke”.
+Zero-shot is the simplest type of prompt, where we directly assign a task to the model.
 Few-shot gives the model a few demonstrations of the task.
@@
 Chain-of-thought (CoT) (Kojima et al., 2022) prompts LLMs to generate the reasoning path (i.e., step-by-step thoughts) for addressing complicated problems.
-For example, “Q: I had 10 apples.
-I gave my neighbor 3 apples.
-Then I bought 4 apples and ate 2.
-How many apples do I have now?
-A: Let’s think step by step:”.
 Based on CoT, self-consistency (Wang et al., 2022b) aggregates a diverse set of reasoning paths to get the most consistent answer.
-Recently, tree-of-thought (ToT) (Yao et al., 2023) further builds a tree of reasoning paths and applies tree search methods to obtain the best path.
+Tree-of-thought (ToT) (Yao et al., 2023) further builds a tree of reasoning paths and applies tree search methods to obtain the best path.
 Self-refine (Madaan et al., 2023) prompts an LLM to provide feedback for its own outputs and refine itself.
 We can even prompt an LLM to write prompts.
-FMs start to demonstrate potential generalization power in the medical field using ICL techniques (Roy et al., 2023; Lyu et al., 2023).
+ICL can provide grounded knowledge to compensate for the lag in pre-training data of FMs and the lack of domain-specific knowledge.
+Retrieval-augmented generation (RAG) is a technique that leverages external knowledge retireval mechanisms to gather additional relevant information for ICL, enhancing the generation quality of FMs (Lewis et al., 2020; Zhao et al., 2024).
+In RAG, FMs (typically, LLMs) first generate queries based on user input, which are then used to retrieve information from a knowledge base or external documents.
+The retrieved context is integrated with the input, allowing FMs to produce more informed and accurate outputs.
+FMs start to demonstrate generalization power in the medical field using ICL techniques (Roy et al., 2023; Lyu et al., 2023).
 For instance, ChatGPT can effectively translate radiology reports into plain language by using well-designed prompts that inform the model about the structure of the report and contents of each paragraph (Lyu et al., 2023).
 Also, the general-purpose SAM performs well on abdominal CT organ segmentation when provided point and bounding box prompts in an oracle manner (Roy et al., 2023).
+Besides, RAG methods show great potential in enhancing the capabilities of LLMs for medical-related NLP tasks (Zhu et al., 2024; Wu et al., 2024).
 3. Foundation Models in Healthcare
@@
 The experiments show that by using a clearer and more structured prompt, the overall translation quality can be increased.
-Roy et al.
-(Roy et al., 2023) show that SAM (Kirillov et al., 2023) has a great generalization ability in abdominal CT organ segmentation, through its point/bounding box prompting.
 Deng et al.
@@
 Chen et al.
-propose Diagnosis of Thought (DoT) prompting (Chen et al., 2023b) to assist professionals with cognitive distortion detection.
+propose Diagnosis of Thought (DoT) prompting (Chen et al., 2023d) to assist professionals with cognitive distortion detection.
 DoT diagnoses mental illness by prompting LLMs to sequentially perform subjectivity assessment, contrastive reasoning, and schema analysis.
+Besides, Zhu et al.
+(Zhu et al., 2024) augment multi-modal electronic health records (EHR) data with a medical knowledge graph via retrieval-augmented generation (RAG).
 3.2. Pre-Training Healthcare Foundation Models
@@
 For instance, SciBERT (Beltagy et al., 2019) and PubMedBERT (Gu et al., 2021) are pre-trained on multi-domain scientific publications and biomedical domain-specific corpora respectively, based on BERT strategy.
-BioLinkBERT (Yasunaga et al., 2022) utilizes links between biomedical documents and is pre-trained with masked language modeling and document relation prediction tasks.
 BioGPT (Luo et al., 2022) is pre-trained on PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/ abstracts following GPT-2 (Radford et al., 2019) for generative language tasks.
@@
 For example, REMEDIS (Azizi et al., 2023) is a medical vision model pre-trained via contrastive learning to extract representative visual features for medical images.
-Vision-language models such as MedCLIP (Wang et al., 2022c), MI-Zero (Lu et al., 2023b), and PLIP (Huang et al., 2023a), are contrastively pre-trained on domain-specific image-text pairs.
+Pai et al.
+develop an FM for cancer imaging biomarker discovery by contrastively training a convolutional vision encoder (Pai et al., 2024).
+Vision-language models such as MedCLIP (Wang et al., 2022c), MI-Zero (Lu et al., 2023c), and PLIP (Huang et al., 2023a), are contrastively pre-trained on domain-specific image-text pairs.
 They achieve positive performance on zero-shot image classification tasks in radiology and pathology.
@@
 For example, BioBERT (Lee et al., 2020) is pre-trained on PubMed abstracts and PubMed Central (PMC [^2] https://www.ncbi.nlm.nih.gov/pmc/) full-text articles with BERT initialization and demonstrates better performance compared with BERT and previous state-of-the-art models when fine-tuned on three biomedical language tasks.
-Similarly, BlueBERT (Peng et al., 2019) continues BERT pre-training on PubMed abstracts and MIMIC-III clinical notes (Johnson et al., 2016).
-Alsentzer et al.
-(Alsentzer et al., 2019) build clinical text models initialized from BERT and BioBERT, and show a significant improvement on five clinical language tasks.
 PMC-LLaMA (Wu et al., 2023d) is an open-source language model fine-tuned from LLaMA (Touvron et al., 2023a) on biomedical academic papers.
@@
 (Singhal et al., 2022) apply instruction tuning on Flan-PaLM (Chung et al., 2022) to obtain Med-PaLM model, achieving state-of-the-art performance on a broad range of medical question answering (MedQA) tasks.
+MMedLM (Qiu et al., 2024) is a multi-lingual language model for medicine, further pre-trained from InternLM (Cai et al., 2024) using over 25B medical-related tokens across six languages.
 PubMedCLIP (Eslami et al., 2021), a CLIP model fine-tuned on medical image-text pairs from PubMed articles, shows outstanding performance on medical visual question answering (MedVQA) tasks.
@@
 Standard fusion methods can be conceptually categorized into three types (Stahlschmidt et al., 2022; Huang et al., 2020): early fusion, joint fusion, and late fusion.
-In early fusion, data from different modalities are combined at the input and passed to the subsequent network (Park et al., 2020; Xie et al., 2019; Zhao et al., 2021; Suresh et al., 2017; Bichindaritz et al., 2021; Fu et al., 2020).
+In early fusion, data from different modalities are combined at the input and passed to the subsequent network (Park et al., 2020; Zhao et al., 2021; Bichindaritz et al., 2021).
 Joint fusion processes data from each modality through independent networks before fusing their representative feature maps as the input of the subsequent network (Braman et al., 2021; Spasov et al., 2018; Kawahara et al., 2018).
@@
 These models extract image and text features with separated networks and project both features into a shared latent space, enabling a joint training network.
-These models have shown great zero-shot transferability when testing on unseen image-text combinations, including medical images and healthcare reports (Liu et al., 2023f).
+These models have shown great zero-shot transferability when testing on unseen image-text combinations, including medical images and healthcare reports (Liu et al., 2023g).
 In addition, GLIP (Li et al., 2022), CoCa (Yu et al., 2022), and FLAVA (Singh et al., 2022) introduce more advanced fusion modules, such as cross-attention and generative models, to better explore cross-modality interaction during pre-training.
@@
 BioViL (Boecking et al., 2022) trains a visual model together with an in-domain (CXR) language model by optimizing contrastive loss and masked language modeling.
-In addition, PubMedCLIP (Eslami et al., 2021), CheXzero (Tiu et al., 2022), and PLIP (Huang et al., 2023a) extend the CLIP model to image-text data in biomedicine, radiology, and pathology respectively.
+PubMedCLIP (Eslami et al., 2021), CheXzero (Tiu et al., 2022), PLIP (Huang et al., 2023a), and EchoCLIP (Christensen et al., 2024) extend the CLIP model to image-text data in biomedicine, radiology, pathology, and echocardiography respectively.
+In addition, PRISM (Shaikovski et al., 2024) is trained on thousands of clinical reports, each grouped with one or more whole slide images (WSIs), and can generate reports based on WSIs and text prompts.
 BiomedJourney (Gu et al., 2023) is a latent diffusion model (Rombach et al., 2022) trained on tuples of pre- and post-disease progression images from patients coupled with a GPT-generated natural language description of the progression, enabling the generation of counterfactual medical images.
@@
 These combined multi-modal inputs are then fused through the Transformer blocks in the LLM, exchanging information via attention layers (Wang et al., 2023a; Liu et al., 2023a, b).
-For instance, Frozen (Tsimpoukelli et al., 2021) concatenates image features and text tokens as the input of a pre-trained generative language model.
-This design enables the vision encoder to update based on information from both modalities.
-Similarly, SimVLM (Wang et al., 2021) fuses image features and text tokens with BERT-like bi-directional attention layers and trains the whole model together.
+For instance, SimVLM (Wang et al., 2021) fuses image features and text tokens with BERT-like bi-directional attention layers and trains the whole model together.
 Flamingo (Alayrac et al., 2022) takes as input visual data interleaved with text and inserts gated xattn-dense layers into the language model to condition it on visual inputs.
@@
 In addition, LLaVA (Liu et al., 2023b) feeds image features and language instructions into LLaMA to train a multi-modal chatbot.
-NExT-GPT (Wu et al., 2023a) is a unified agent understanding multiple modalities, including text, images, videos, and audio.
+Besides vision and language, recent large multi-modal models (LMMs) are extending to more modalities.
+For example, NExT-GPT (Wu et al., 2023a) is a unified agent understanding multiple modalities, including text, image, video, and audio.
+GPT-4o (OpenAI, 2024), which is an optimized version of GPT-4, works with any combination of text, audio, image, and video quickly, enabling more natural human-computer interaction.
 Since LLMs show great potential in the medical field (Nori et al., 2023), these methods are applied to medical data fusion as well.
@@
 Surprisingly, GPT-4 generates high-quality visual question-answering conversations even if it has access only to the text.
+CheXagent (Chen et al., 2024) bridges a vision encoder with a clinical LLM by adopting a Q-Former (Li et al., 2023c), and trains with curated instruction data.
 These models show outstanding zero-shot/few-shot performance on downstream medical tasks that require multi-modality understanding.
 Current multi-modal FMs predominantly focus on language and vision, given their ubiquitous use as the most common modalities.
-For FM applications in the healthcare field, there are only handful examples including genomics (Ding et al., 2023b), electrocardiogram (Chang et al., 2023a), electroencephalogram (Cui et al., 2023a), and tabular data (Wang et al., 2023c).
+For FM applications in the healthcare field, there are only a handful of examples including molecules (Liu et al., 2023c), genomics (Ding et al., 2023b), electrocardiogram (Chang et al., 2023a), electroencephalogram (Cui et al., 2023a), and tabular data (Wang et al., 2023c).
 We believe that joint-modal pre-training techniques, combined with LLMs’ logical reasoning capabilities, could benefit healthcare data fusion of modalities beyond the scope of language and vision.
@@
 Towards applying FMs in the healthcare system safely and responsibly, we must address a variety of data quantity hurdles.
-The trade-off of data quantity is between the limited information provided by a constrained dataset and the essential information required for training a robust model on the specific healthcare task.
+The trade-off of data quantity lies between the limited information provided by a constrained dataset and the essential information required for training a robust model on the specific healthcare task.
 Such a disparity can result in inadequate downstream model training, yielding non-robust, inaccurate, and even unreliable model outcomes.
@@
 In addition, we must deal with the costly real-world dataset curation processes including data collection, cleaning, and annotation.
-While deep learning has shown its promise, the barrier of limited data access in healthcare continues to be a significant roadblock for scalable AI-based research.
+While deep learning has shown its promise, the barrier of limited data access in healthcare remains a significant roadblock for scalable AI-based research.
 The emergence of FMs holds the potential to alleviate the data quantity challenges in healthcare applications.
@@
 (Wu et al., 2023c) demonstrate that general large language models (LLMs) such as ChatGPT (OpenAI, 2021) and GPT-4 (OpenAI, 2023) exhibit strong capabilities in handling radiology natural language inference (NLI) tasks even with limited data.
+General FMs can efficiently retrieve external information from public data sources for downstream tasks using retrieval-augmented generation (RAG) techniques.
+RAG is a highly effective method for enabling FMs to acquire grounded domain knowledge which is publicly accessible but not included in model pre-training.
+For instance, BioRAG (Wang et al., 2024) leverages an LLM to adaptively select knowledge sources from papers, search engines, and biological data hubs for biological question-reasoning.
+REALM (Zhu et al., 2024) extracts comprehensive representations for EHR data by fusing EHR with information retrieved from a medical knowledge graph, without training on task-specific data.
 In addition to general FMs, leveraging healthcare-focused FMs represents a reasonable direction for pursuing data efficiency, significantly reducing the data requirement in downstream healthcare applications.
@@
 (Gilson et al., 2022) reveal that ChatGPT (OpenAI, 2021) achieves comparable human-level performance on medical question-answering tasks, showcasing the potential for LLMs to enhance the efficiency of text annotation in a more flexible manner.
-Med-PaLM 2 (Singhal et al., 2023), an LLM fine-tuned with healthcare data, has performed comparably to human clinical professionals on medical knowledge retrieval, reasoning, and question answering.
+Med-PaLM 2 (Singhal et al., 2023a), an LLM fine-tuned with healthcare data, has performed comparably to human clinical professionals on medical knowledge retrieval, reasoning, and question answering.
 This suggests that integrating Med-PaLM 2 into medical data annotation workflows could lead to more accurate and efficient annotation processes, harnessing the power of LLMs in the medical domain.
 While LLMs have showcased the ability to produce grammatically accurate and human-like text (Thirunavukarasu et al., 2023; Gilson et al., 2022), they still encounter challenges in specialized healthcare contexts.
-A noticeable knowledge gap in healthcare text still remains between general-domain FMs and medicine professionals, resulting in inaccurate responses to realistic patient inquiries, such as in cardiovascular disease prevention (Sarraju et al., 2023).
+A noticeable knowledge gap in healthcare text remains between general-domain FMs and medicine professionals, resulting in inaccurate responses to realistic patient inquiries, such as in cardiovascular disease prevention (Sarraju et al., 2023).
 Besides, Liao et al.
 (Liao et al., 2023) also highlight that the linguistic attributes of medical text generated by ChatGPT considerably differ from those of human experts.
-Advancements are required to improve the utilization of text-based FMs for improving the performance of medical text annotation.
+Advancements are required to improve the utilization of text-based FMs for more accurate and efficient medical text annotation.
 PathAsst (Sun et al., 2023) leverages FMs as a generative AI assistant to transform predictive analytics in pathology.
@@
 Also, Liu et al.
-(Liu et al., 2023e) develop DeID-GPT as a de-identification framework powered by GPT-4 (OpenAI, 2023), designed to detect and eliminate identification information from clinical text records automatically.
-In contrast to conventional methods, DeID-GPT utilizes in-context learning, which requires no code or procedural changes when applying the solution to different hospitals, different languages, and different data formats.
-It improves reliability by concealing sensitive information while preserving the original structure of the clinical text.
+(Liu et al., 2023f) develop DeID-GPT as a de-identification framework powered by GPT-4 (OpenAI, 2023), designed to detect and eliminate identification information from clinical text records automatically.
+Similarly, Wu et al.
+(Wu et al., 2024) construct a medical knowledge graph by prompting an LLM to identify and extract entities from segmented medical documents.
+The LLM is further employed to detect relevance and link the entities, forming a three-level hierarchy.
 Moreover, Tang et al.
-(Tang et al., 2023b) utilize the language model BART (Lewis et al., 2019) to enhance medical dialogue generation by focusing on domain-specific terminology.
+(Tang et al., 2023b) utilize a language model to enhance medical dialogue generation by focusing on domain-specific terminology.
 Empirical results highlight the model’s efficiency in generating responses based on the historical context of dialogue exchanges between medical professionals and patients.
@@
 Such a task can also involve annotating regions of interest in histopathological slides to identify cancerous cells or specific tissue types of clinical significance.
-Medical image annotation plays a pivotal role in modern healthcare to help extract and interpret valuable information from complex medical image examinations.
-This process is essential for both clinical diagnosis and research purposes, as healthcare professionals require computer-aided algorithms to efficiently identify, search, manage, and interpret disease indicators (Yin and Bi, 2018).
-Robust segmentation models can be transformed into powerful tools for healthcare image annotation, thus leading to a substantial reduction in annotation costs and the facilitation of large-scale dataset curation spanning a wider range of healthcare imaging applications (Gao et al., 2023).
+Medical image annotation is essential in modern healthcare as it helps to extract and interpret valuable information from complex medical image examinations.
+This is necessary for clinical diagnosis and research, as healthcare professionals require computer-aided algorithms to efficiently identify, search, manage, and interpret disease indicators (Yin and Bi, 2018).
+Robust segmentation models can be transformed into powerful tools for healthcare image annotation, leading to a substantial reduction in annotation costs and the facilitation of large-scale dataset curation spanning a wider range of healthcare imaging applications (Gao et al., 2023).
 For instance, Qu et al.
@@
 The outcome consistency and reliability of directly using SAM in medical image segmentation still need to be improved by properly refining FMs (Zhang and Jiao, 2023).
-Several studies have focused on adapting FMs to the complexities of medical data, aiming to improve the segmentation performance, which potentially benefits the image annotation performance.
+Several studies have focused on adapting FMs to the complexities of medical data, aiming to improve segmentation performance, which potentially benefits image annotation.
 Zhang et al.
@@
 (Hu et al., 2023b) enhance SAM’s performance through efficient modifications by refining the lightweight task-specific prediction head while keeping the SAM encoder unaltered.
-Empirical validations indicate that finetuning SAM can significantly elevate its performance on medical image datasets.
+Empirical validations indicate that finetuning SAM can significantly elevate its performance with medical images.
 Technical challenges remain in adapting SAM to the medical domain efficiently and scalably.
@@
 In particular, SAM pre-trained in general vision lacks the specialized understanding of medical anatomy, pathology, and radiology necessary for precise image analysis.
-Data- or parameter-efficient fine-tuning is effective while it is not scalable.
+Data- or parameter-efficient fine-tuning is effective while not scalable.
 Moving forward, more research efforts are required towards a more scalable medical knowledge transfer pattern for better utilizing and adapting SAM or other general-purpose FMs in medical image and text annotation.
@@
 Safeguarding healthcare data privacy is crucial given their critical roles in clinical research, diagnosis, treatment, and disease prevention (Gkoulalas-Divanis et al., 2014).
-Laws and regulations have been enacted to ensure the protection of these highly sensitive personal records (Clayton et al., 2023; Malin et al., 2010), fostering the trust between patients and medical institutions.
-Traditionally, healthcare data is well protected to preserve privacy by strategies such as cryptographic methods and access control mechanisms (Gertz and Jajodia, 2007; Aggarwal and Yu, 2008).
-Nevertheless, these approaches encounter limitations in providing absolute patient anonymity, whereby the private and confidential information of patients remains undisclosed, particularly when disseminating patient data (Fung et al., 2010).
+Laws and regulations have been enacted to safeguard these highly sensitive personal records (Clayton et al., 2023; Malin et al., 2010), which helps to build trust between patients and medical institutions.
+Traditionally, healthcare data is well protected to preserve privacy through techniques like encryption and access control mechanisms (Gertz and Jajodia, 2007; Aggarwal and Yu, 2008).
+Nevertheless, these approaches encounter limitations in providing absolute patient anonymity, whereby private and confidential information of patients remains undisclosed, particularly when disseminating patient data (Fung et al., 2010).
 The challenge arises due to the necessity of distributing data to a broad and potentially unidentified group of recipients.
 To solve this, healthcare data is typically altered to remove private information before being used (de Wolf, 2012; Willenborg and De Waal, 2012; Kosorok and Laber, 2019).
-Perturbation methods are used to add controlled noise or randomness to data to protect privacy while allowing data analysis (Turgay et al., 2023).
+Perturbation methods involve adding controlled noise or randomness to data to protect privacy while allowing data analysis (Turgay et al., 2023).
 These methods are leveraged at a reduced level of intensity to mitigate the potential deterioration in the integrity of subsequent secondary data analyses (Reiter, 2012).
@@
 The advent of data-centric FMs introduces both opportunities and challenges in healthcare data privacy.
-On one hand, FMs offer improved capabilities for data generation, providing unidentified synthetic data (Tang et al., 2023a; Shibata et al., 2023).
+On the one hand, FMs offer improved capabilities for data generation, providing unidentified synthetic data (Tang et al., 2023a; Shibata et al., 2023).
 On the other hand, FMs tend to memorize their training data due to their massive parameters, resulting in the generation of similar outputs that potentially contain private patient information (Zhou et al., 2023b; Carlini et al., 2021).
@@
 The growing use of synthetic data represents a solution to protect private information that is in line with the key promise of FMs (Hu et al., 2023a).
-Deep generative models can now produce high-quality synthetic medical data with similar characteristics to the original data but without identical information (Naresh et al., 2023), thus preserving the functionality while avoiding risks of privacy leakage.
+Deep generative models can now produce high-quality synthetic medical data with similar characteristics to the original data but without identical information (Naresh et al., 2023), thus preserving functionality while avoiding privacy leakage.
 For instance, recent advances in FM-based data generation methods have enabled the anonymization of medical images and texts, supporting data privacy while maintaining a wide utility (Dong et al., 2022b).
@@
 Although FMs shed light on addressing medical data privacy issues, we have recognized that FMs themselves have brought up emerging privacy protection challenges.
-Numerous studies have extensively investigated FMs of their propensity to memorize training data across different domains, a phenomenon referred to as memorization (Carlini et al., 2019; Thakkar et al., 2021; Zhang et al., 2021b; Carlini et al., 2022; Kandpal et al., 2022; Lee et al., 2021).
+Numerous studies have extensively investigated FMs of their propensity to memorize training data across different domains, a phenomenon referred to as memorization (Carlini et al., 2019, 2022; Kandpal et al., 2022; Lee et al., 2021).
 FMs have been trained on private datasets (Carlini et al., 2023).
@@
 The Biomedical Language Understanding & Reasoning Benchmark (BLURB) (Gu et al., 2021) focuses on biomedical datasets and improves BLUE by including question-answering tasks which are frequently employed to evaluate large language models (LLMs).
+Medbench (Liu et al., 2024a) provides over 300K Chinese questions covering 43 clinical specialties, an automatic evaluation system, and dynamic evaluation mechanisms for unbiased and reliable assessment of Chinese medical LLMs.
 Table 1 displays a representative list of public medical databases, benchmarks, and datasets that can facilitate FM training, inference, and evaluation.
 Table 1. Public medical databases, benchmarks, and datasets that can facilitate foundation model development and application in medicine.
-Database/Benchmark/Dataset Description Task type MIMIC-III (Johnson et al., 2016) Critical care data for over 40,000 patients Text BLUE (Peng et al., 2019) 5 language tasks with 10 biomedical and clinical text datasets Text MedMentions (Mohan and Li, 2019) 4,392 papers annotated by experts with mentions of UMLS entities Text webMedQA (He et al., 2019) 63,284 real-world Chinese medical questions with 300K answers Text PubMedQA (Jin et al., 2019) 1K expert-annotated, 61.2K unlabeled, and 211.3K artificially generated biomedical QA instances Text BLURB (Gu et al., 2021) 13 biomedical NLP datasets in 6 tasks Text CBLUE (Zhang et al., 2021a) A Chinese medical NLP benchmark with 18 datasets Text MedQA-USMLE (Jin et al., 2021) 61,097 multiple choice questions based on USMLE in three languages Text MedMCQA (Pal et al., 2022) 194K multiple-choice questions covering 2.4K healthcare topics Text MultiMedQA (Singhal et al., 2022) 6 existing and 1 online-collected medical QA dataset Text ClinicalTrials.gov [^3] https://clinicaltrials.gov/ An online database of clinical research studies, including clinical trials and observational studies Text Medical Meadow (Han et al., 2023) 16M medical QA pairs collected from 9 sources Text Huatuo-26M (Li et al., 2023d) 26M Chinese medical QA pairs Text ChestXray-NIHCC (Wang et al., 2017) 100K radiographs with labels from more than 30,000 patients Imaging DeepLesion (Yan et al., 2018) 32K CT scans with annotations and semantic labels from radiological reports Imaging Kather Colon Dataset (Kather et al., 2018) 100K histological images of human colorectal cancer and healthy tissue Imaging CheXpert (Irvin et al., 2019) 224,316 chest radiographs of 65,240 patients Imaging EchoNet-Dynamic (Ouyang et al., 2020) 10,030 expert-annotated echocardiogram videos Imaging ISIC [^4] https://www.isic-archive.com/ An archive containing 23K skin lesion images with labels Imaging Med-MNIST v2 (Yang et al., 2023a) 12 2D and 6 3D datasets for biomedical image classification Imaging AbdomenAtlas-8K (Qu et al., 2023) 8,448 CT volumes with per-voxel annotated eight abdominal organs Imaging Virchow (Vorontsov et al., 2023) 1.5M pathological scans from 120K patients Imaging RETFound (Zhou et al., 2023a) Unannotated retinal images, containing 904,170 CFPs and 736,442 OCT scans Imaging dbSNP (Sherry et al., 2001) A collection of human single nucleotide variations, microsatellites, and small-scale insertions and deletions Genomics ENCODE (Consortium et al., 2012) A platform of genomics data and encyclopedia with integrative-level and ground-level annotations Genomics 1000 Genomes Project (Consortium et al., 2015) A comprehensive catalog of human genetic variations Genomics ChEMBL (Gaulton et al., 2012) 20M bioactivity measurements for 2.4M distinct compounds and 15K protein targets Drug DrugBank (Wishart et al., 2018) A web-enabled structured database of molecular information about drugs Drug PubChem (Kim et al., 2023) A collection of 900+ sources of chemical information data Drug TCGA [^5] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A landmark cancer genomics program, molecularly characterized over 20,000 primary cancer and matched normal samples spanning 33 cancer types Multi-modal MIMIC-CXR (Johnson et al., 2019) 227,835 chest imaging studies with free-text reports for 65,379 patients Multi-modal MIMIC-IV (Johnson et al., 2020) Clinical information for hospital stays of over 60,000 patients Multi-modal SwissProtCLAP (Liu et al., 2023g) 441K text-protein sequence pairs Multi-modal PMC-VQA (Zhang et al., 2023d) 227K VQA pairs of 149K images of various modalities or diseases Multi-modal MedMD (Wu et al., 2023e) 15.5M 2D scans and 180k 3D radiology scans with textual descriptions Multi-modal PathCap (Sun et al., 2023) 142K pathology image-caption pairs from various sources Multi-modal PathInstruct (Sun et al., 2023) 180K samples of LLM-generated instruction-following data Multi-modal OpenPath (Huang et al., 2023a) 208,414 pathology images paired with natural language descriptions Multi-modal Quilt-1M (Ikezogwo et al., 2023) 1M image-text pairs for histopathology Multi-modal PMC-OA (Lin et al., 2023b) 1.6M fine-grained biomedical image-text pairs Multi-modal
+Database/Benchmark/Dataset Description Task type ClinicalTrials.gov [^3] https://clinicaltrials.gov/ An online database of clinical research studies, including clinical trials and observational studies Text MIMIC-III (Johnson et al., 2016) Critical care data for over 40,000 patients Text BLUE (Peng et al., 2019) 5 language tasks with 10 biomedical and clinical text datasets Text MedMentions (Mohan and Li, 2019) 4,392 papers annotated by experts with mentions of UMLS entities Text webMedQA (He et al., 2019) 63,284 real-world Chinese medical questions with 300K answers Text PubMedQA (Jin et al., 2019) 1K expert-annotated, 61.2K unlabeled, and 211.3K artificially generated biomedical QA instances Text BLURB (Gu et al., 2021) 13 biomedical NLP datasets in 6 tasks Text CBLUE (Zhang et al., 2021) A Chinese medical NLP benchmark with 18 datasets Text MedQA-USMLE (Jin et al., 2021) 61,097 multiple choice questions based on USMLE in three languages Text MedMCQA (Pal et al., 2022) 194K multiple-choice questions covering 2.4K healthcare topics Text MultiMedQA (Singhal et al., 2022) 6 existing and 1 online-collected medical QA dataset Text Medical Meadow (Han et al., 2023) 16M medical QA pairs collected from 9 sources Text Huatuo-26M (Li et al., 2023d) 26M Chinese medical QA pairs Text GAP-Replay (Chen et al., 2023a) 48.1B tokens from 4 medical corpora including guidelines, abstracts, papers, and replay Text BiMed1.3M (Pieri et al., 2024) An English and Arabic bilingual dataset of 1.3M medical QA and chat samples Text MMedC (Qiu et al., 2024) A multilingual medical corpus containing over 25.5B tokens Text MedBench (Liu et al., 2024a) 300,901 Chinese questions covering 43 clinical specialties, combined with an automatic evaluation system Text ISIC [^4] https://www.isic-archive.com/ An archive containing 23K skin lesion images with labels Imaging ChestXray-NIHCC (Wang et al., 2017) 100K radiographs with labels from more than 30,000 patients Imaging DeepLesion (Yan et al., 2018) 32K CT scans with annotations and semantic labels from radiological reports Imaging Kather Colon Dataset (Kather et al., 2018) 100K histological images of human colorectal cancer and healthy tissue Imaging CheXpert (Irvin et al., 2019) 224,316 chest radiographs of 65,240 patients Imaging EchoNet-Dynamic (Ouyang et al., 2020) 10,030 expert-annotated echocardiogram videos Imaging Med-MNIST v2 (Yang et al., 2023a) 12 2D and 6 3D datasets for biomedical image classification Imaging AbdomenAtlas-8K (Qu et al., 2023) 8,448 CT volumes with per-voxel annotated eight abdominal organs Imaging Virchow (Vorontsov et al., 2023) 1.5M pathological scans from 120K patients Imaging AbdomenAtlas-8K 8,448 CT volumes with per-voxel annotated eight abdominal organs Imaging RETFound (Zhou et al., 2023a) Unannotated retinal images, containing 904,170 CFPs and 736,442 OCT scans Imaging dbSNP (Sherry et al., 2001) A collection of human single nucleotide variations, microsatellites, and small-scale insertions and deletions Genomics ENCODE (Consortium et al., 2012) A platform of genomics data and encyclopedia with integrative-level and ground-level annotations Genomics 1000 Genomes Project (Consortium et al., 2015) A comprehensive catalog of human genetic variations Genomics ChEMBL (Gaulton et al., 2012) 20M bioactivity measurements for 2.4M distinct compounds and 15K protein targets Drug DrugBank (Wishart et al., 2018) A web-enabled structured database of molecular information about drugs Drug PubChem (Kim et al., 2023) A collection of 900+ sources of chemical information data Drug DrugChat (Liang et al., 2023) 143,517 question-answer pairs covering 10,834 drug compounds, collected from PubChem and ChEMBL Drug TCGA [^5] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A landmark cancer genomics program, molecularly characterized over 20,000 primary cancer and matched normal samples spanning 33 cancer types Multi-modal MIMIC-CXR (Johnson et al., 2019) 227,835 chest imaging studies with free-text reports for 65,379 patients Multi-modal MIMIC-IV (Johnson et al., 2020) Clinical information for hospital stays of over 60,000 patients Multi-modal SwissProtCLAP (Liu et al., 2023h) 441K text-protein sequence pairs Multi-modal PMC-VQA (Zhang et al., 2023d) 227K VQA pairs of 149K images of various modalities or diseases Multi-modal MedMD (Wu et al., 2023e) 15.5M 2D scans and 180k 3D radiology scans with textual descriptions Multi-modal PathCap (Sun et al., 2023) 142K pathology image-caption pairs from various sources Multi-modal PMC-OA (Lin et al., 2023b) 1.6M fine-grained biomedical image-text pairs Multi-modal PathInstruct (Sun et al., 2023) 180K samples of LLM-generated instruction-following data Multi-modal Quilt-1M (Ikezogwo et al., 2023) 1M image-text pairs for histopathology Multi-modal OpenPath (Huang et al., 2023a) 208,414 pathology images paired with natural language descriptions Multi-modal Chi-Med-VL (Liu et al., 2023d) 580,014 image-text pairs and 469,441 question-answer pairs for general healthcare in Chinese Multi-modal SAT-DS (Zhao et al., 2023) 11,462 scans with 142,254 segmentation annotations spanning 8 human body regions Multi-modal
 Current benchmarks are largely focused on static datasets, which are unable to capture the full complexities of how FMs behave in dynamic environments with continuous variants, real-time adjustments, and interaction with humans.
@@
 The lack of such alignment could pose misleading estimates of FM capabilities leading to harmful results.
-In fact, LLM alignment evaluation from perspectives of reliability, safety, fairness, explainability, and robustness underscores the importance of LLM trustworthiness (Liu et al., 2023d).
+In fact, LLM alignment evaluation from perspectives of reliability, safety, fairness, explainability, and robustness underscores the importance of LLM trustworthiness (Liu et al., 2023e).
 Yet we lack substantial research benchmarks for FM trustworthiness evaluation in the healthcare field.
@@
 For instance, Chen et al.
-(Chen et al., 2023c) explore reference-free evaluation methods that prompt ChatGPT to score the quality of model-generated texts without a pre-defined ground truth.
+(Chen et al., 2023e) explore reference-free evaluation methods that prompt ChatGPT to score the quality of model-generated texts without a pre-defined ground truth.
 Chiang et al.
@@
 In addition, the rigorous audit scheme in the process of dataset curation could alleviate the harmful information that may already exist in the massive training data.
-From a model perspective, the inherent bias in FM architectures should be recognized by healthcare stakeholders and developers.
-Currently, FM architectures and training patterns lack the defense and detection capability against adversarial manipulation.
-FMs could propagate all knowledge learned from training data without any regulation of biased knowledge detection.
+From a model perspective, healthcare stakeholders and developers should acknowledge the inherent bias present in FM architectures.
+Currently, FM architectures and training patterns lack defense and detection capabilities against adversarial manipulation.
+FMs could propagate all knowledge learned from training data without a proper mechanism to identify and address biased knowledge.
 This vulnerability can potentially hurt healthcare stakeholders by generating false or harmful information.
@@
 Popularizing the standardized guideline documents, such as the ”datasheet” (Gebru et al., 2021), is crucial to simplify metadata preparation.
-These efforts are helpful for us to understand the comprehensive details of the dataset, enhancing the wide adoption of clinical data to accelerate the progress of FM research in healthcare.
+These efforts are helpful for us to understand the comprehensive details of the dataset, which in turn promotes the wide adoption of clinical data to advance FM research in healthcare.
 10. Conclusions
@@
 Higher magnifications reveal finer attributes, while lower magnifications facilitate the understanding of overall tissue morphology (Ashtaiwi, 2022).
-Due to the high resolution of histopathological images, current FMs mostly take split image patches as input instead of whole slides (Huang et al., 2023a; Lu et al., 2023a).
+Due to the high resolution of histopathological images, current FMs mostly take split image patches as input instead of whole slides (Huang et al., 2023a; Lu et al., 2023b).
 Molecular information typically encompasses various categories of biomarkers, including DNA, RNA, proteins, metabolites, and more recently microbiome profiles (Mishra and Verma, 2010).
@@
 Table 2 highlights healthcare FMs, including their model structure, initialization, pre-training data, and the link reference to the project.
-Table 2. Typical foundation models in healthcare and medicine.
-Model Base Initial Model Weights Pre-training Data SciBERT (Beltagy et al., 2019) BERT (Devlin et al., 2018) - Semantic Scholar (Ammar et al., 2018) Clinical BERT (Alsentzer et al., 2019) BERT BERT MIMIC-III (Johnson et al., 2016) BlueBERT (Peng et al., 2019) BERT - PubMed [^6] https://pubmed.ncbi.nlm.nih.gov/ + MIMIC-III BioBERT (Lee et al., 2020) BERT BERT PubMed + PMC [^7] https://www.ncbi.nlm.nih.gov/pmc/ PubMedBERT (Gu et al., 2021) BERT - PubMed BioLinkBERT (Yasunaga et al., 2022) BERT - PubMed BioGPT (Luo et al., 2022) GPT-2 (Radford et al., 2019) - PubMed Clinical-T5 (Lehman and Johnson, 2023) T5 (Raffel et al., 2020) T5 MIMIC-III + MIMIC-IV (Johnson et al., 2020) HIPT (Chen et al., 2022) DINO (Caron et al., 2021) - TCGA [^8] https://www.cancer.gov/ccg/research/genome-sequencing/tcga CTransPath (Wang et al., 2022d) SRCL (Wang et al., 2022d) - TCGA + PAIP [^9] http://www.wisepaip.org/paip/ RETFound (Zhou et al., 2023a) MAE (He et al., 2022) Vision Transformer (Dosovitskiy et al., 2020) RETFound REMEDIS (Azizi et al., 2023) SimCLR (Chen et al., 2020) BiT (Beyer et al., 2022) MIMIC-IV + CheXpert (Irvin et al., 2019) Virchow (Vorontsov et al., 2023) DINOv2 (Oquab et al., 2023) - Virchow PubMedCLIP (Eslami et al., 2021) CLIP (Radford et al., 2021) CLIP ROCO (Pelka et al., 2018) CheXzero (Tiu et al., 2022) CLIP CLIP MIMIC-CXR (Johnson et al., 2019) MedCLIP (Wang et al., 2022c) CLIP Clinical BERT + SwinTransformer (Liu et al., 2021) CheXpert + MIMIC-CXR BiomedCLIP (Zhang et al., 2023e) CLIP PubMedBERT PMC-15M (Zhang et al., 2023e) PMC-CLIP (Lin et al., 2023b) CLIP PubMedBERT PMC-OA (Lin et al., 2023b) MI-Zero (Lu et al., 2023b) CLIP HistPathGPT (Lu et al., 2023b) + CTransPath (Wang et al., 2022e) ARCH (Gamper and Rajpoot, 2021) PLIP (Huang et al., 2023a) CLIP CLIP OpenPath (Huang et al., 2023a) QuiltNet (Ikezogwo et al., 2023) CLIP CLIP Quilt-1M (Ikezogwo et al., 2023) KAD (Zhang et al., 2023c) CLIP - MIMIC-CXR + UMLS (Bodenreider, 2004) CONCH (Lu et al., 2023a) CoCa (Yu et al., 2022) - PubMed + PMC Med-Flamingo (Moor et al., 2023b) Flamingo (Alayrac et al., 2022) Flamingo MTB (Moor et al., 2023b) + PMC-OA Med-PaLM (Singhal et al., 2022) PaLM (Chowdhery et al., 2022) PaLM MultiMedQA (Singhal et al., 2022) ChatDoctor (Yunxiang et al., 2023) LLaMA (Touvron et al., 2023a) LLaMA HealthCareMagic-100k (Yunxiang et al., 2023) PMC-LLaMA (Wu et al., 2023d) LLaMA LLaMA MedC (Wu et al., 2023d) Clinical LLaMA-LoRA (Gema et al., 2023) LLaMA LLaMA / PMC-LLaMA MIMIC-IV (Johnson et al., 2020) BioMedGPT (Luo et al., 2023) LLaMA 2 (Touvron et al., 2023b) LLaMA 2 S2ORC (Lo et al., 2019) LLaVA-Med (Li et al., 2023e) LLaVA (Liu et al., 2023b) LLaVA PMC-15M (Zhang et al., 2023e) + GPT-4 (OpenAI, 2023) Qilin-Med-VL (Liu et al., 2023c) LLaVA Chinese-LLaMA2 + CLIP ChiMed-VL (Liu et al., 2023c) MedVInT (Zhang et al., 2023d) - PMC-LLaMA + PMC-CLIP PMC-VQA (Zhang et al., 2023d) RadFM (Wu et al., 2023e) - PMC-LLaMA MedMD (Wu et al., 2023e) ESM-2 (Lin et al., 2023a) Transformer (Vaswani et al., 2017) - UniRef (Suzek et al., 2015) GET (Fu et al., 2023) Transformer - GET AlphaMissense (Cheng et al., 2023a) AlphaFold (Jumper et al., 2021) - PDB (Berman et al., 2000) + UniRef
+Table 2. Typical foundation models in healthcare and medicine. A star (*) after the pre-training data shows that the authors constructed the data with more than two sources.
+Model Base Initial Model Weights Pre-training Data SciBERT (Beltagy et al., 2019) BERT (Devlin et al., 2018) - Semantic Scholar (Ammar et al., 2018) Clinical BERT (Alsentzer et al., 2019) BERT BERT MIMIC-III (Johnson et al., 2016) BlueBERT (Peng et al., 2019) BERT - PubMed [^6] https://pubmed.ncbi.nlm.nih.gov/ + MIMIC-III BioBERT (Lee et al., 2020) BERT BERT PubMed + PMC [^7] https://www.ncbi.nlm.nih.gov/pmc/ PubMedBERT (Gu et al., 2021) BERT - PubMed BioLinkBERT (Yasunaga et al., 2022) BERT - PubMed BioGPT (Luo et al., 2022) GPT-2 (Radford et al., 2019) - PubMed Med-PaLM (Singhal et al., 2022) PaLM (Chowdhery et al., 2022) PaLM MedQA (Jin et al., 2021) Clinical-T5 (Lehman and Johnson, 2023) T5 (Raffel et al., 2020) T5 MIMIC-III + MIMIC-IV (Johnson et al., 2020) ChatDoctor (Yunxiang et al., 2023) LLaMA (Touvron et al., 2023a) LLaMA HealthCareMagic [^8] https://www.askadoctor24x7.com/ PMC-LLaMA (Wu et al., 2023d) LLaMA LLaMA MedC (Wu et al., 2023d) Med-PaLM 2 (Singhal et al., 2023b) PaLM 2 - MedQA Clinical LLaMA-LoRA (Gema et al., 2023) LLaMA LLaMA / PMC-LLaMA MIMIC-IV BioMedGPT (Luo et al., 2023) LLaMA 2 (Touvron et al., 2023b) LLaMA 2 S2ORC (Lo et al., 2019) HIPT (Chen et al., 2022) DINO (Caron et al., 2021) - TCGA [^9] https://www.cancer.gov/ccg/research/genome-sequencing/tcga CTransPath (Wang et al., 2022d) SRCL (Wang et al., 2022d) - TCGA + PAIP [^10] http://www.wisepaip.org/paip/ RETFound (Zhou et al., 2023a) MAE (He et al., 2022) Vision Transformer (Dosovitskiy et al., 2020) * Virchow (Vorontsov et al., 2023) DINOv2 (Oquab et al., 2023) - * REMEDIS (Azizi et al., 2023) SimCLR (Chen et al., 2020) BiT (Beyer et al., 2022) MIMIC-IV + CheXpert (Irvin et al., 2019) UNI (Chen et al., 2023b) DINOv2 - Mass-100K* (Chen et al., 2023b) RudolfV (Dippel et al., 2024) DINOv2 DINOv2 * Pai et al.
+(Pai et al., 2024) SimCLR - * Prov-GigaPath (Xu et al., 2024) MAE - Prov-Path* (Xu et al., 2024) PubMedCLIP (Eslami et al., 2021) CLIP (Radford et al., 2021) CLIP ROCO (Pelka et al., 2018) CheXzero (Tiu et al., 2022) CLIP CLIP MIMIC-CXR (Johnson et al., 2019) MedCLIP (Wang et al., 2022c) CLIP Clinical BERT + SwinTransformer (Liu et al., 2021) CheXpert + MIMIC-CXR BiomedCLIP (Zhang et al., 2023e) CLIP PubMedBERT PMC-15M* (Zhang et al., 2023e) PMC-CLIP (Lin et al., 2023b) CLIP PubMedBERT PMC-OA* (Lin et al., 2023b) MedVInT (Zhang et al., 2023d) - PMC-LLaMA + PMC-CLIP PMC-VQA* (Zhang et al., 2023d) LLaVA-Med (Li et al., 2023e) LLaVA (Liu et al., 2023b) LLaVA PMC-15M (Zhang et al., 2023e) + GPT-4 (OpenAI, 2023) MI-Zero (Lu et al., 2023c) CLIP HistPathGPT (Lu et al., 2023c) + CTransPath (Wang et al., 2022e) ARCH (Gamper and Rajpoot, 2021) PLIP (Huang et al., 2023a) CLIP CLIP OpenPath* (Huang et al., 2023a) QuiltNet (Ikezogwo et al., 2023) CLIP CLIP Quilt-1M* (Ikezogwo et al., 2023) CONCH (Lu et al., 2023b) CoCa (Yu et al., 2022) - PubMed + PMC Med-Flamingo (Moor et al., 2023b) Flamingo (Alayrac et al., 2022) Flamingo MTB (Moor et al., 2023b) + PMC-OA KAD (Zhang et al., 2023c) CLIP - MIMIC-CXR + UMLS (Bodenreider, 2004) RadFM (Wu et al., 2023e) - PMC-LLaMA MedMD* (Wu et al., 2023e) Qilin-Med-VL (Liu et al., 2023d) LLaVA Chinese-LLaMA2 + CLIP ChiMed-VL* (Liu et al., 2023d) PathChat (Lu et al., 2023a) LLaVA LLaMA 2 + UNI PathChatInstruct* (Lu et al., 2023a) CheXagent (Chen et al., 2024) BLIP-2 (Li et al., 2023c) Mistral 7B (Jiang et al., 2023) CheXinstruct* (Chen et al., 2024) EchoCLIP (Christensen et al., 2024) CLIP CLIP * PRISM (Shaikovski et al., 2024) CoCa BioGPT + Virchow * Med-Gemini (Saab et al., 2024) Gemini (Reid et al., 2024) Gemini * RadFound (Liu et al., 2024b) - - RadVLCorpus* (Liu et al., 2024b) ESM-2 (Lin et al., 2023a) Transformer (Vaswani et al., 2017) - UniRef (Suzek et al., 2015) GET (Fu et al., 2023) Transformer - * AlphaMissense (Cheng et al., 2023a) AlphaFold (Jumper et al., 2021) - PDB (Berman et al., 2000) + UniRef MoleculeSTM (Liu et al., 2023c) CLIP - PubChem (Kim et al., 2023)